### v1

# Poisoning with A Pill: Circumventing Detection in Federated Learning

Abstract
Without direct access to the client’s data, federated learning (FL) is well-known for its unique strength in data privacy protection among existing distributed machine learning techniques. However, its distributive and iterative nature makes FL inherently vulnerable to various poisoning attacks, including model poisoning attacks and data poisoning attacks. To counteract these threats, extensive defenses have been proposed to filter out malicious clients, using various detection metrics. Based on our analysis of existing attacks and defenses, we find that there is a lack of attention to model redundancy. In neural networks, various model parameters contribute differently to the model’s performance. However, existing attacks in FL manipulate all the model update parameters with the same strategy, making them easily detectable by common defenses. Meanwhile, the defenses also tend to analyze the overall statistical features of the entire model updates, leaving room for sophisticated attacks. Based on these observations, this paper proposes a generic and attack-agnostic augmentation approach designed to enhance the effectiveness and stealthiness of existing FL poisoning attacks against detection in FL, pointing out the inherent flaws of existing defenses and exposing the necessity of fine-grained FL security. Specifically, we employ a three-stage methodology that strategically constructs, generates, and injects poison (generated by existing attacks) into a pill (a tiny subnet with a novel structure) during the FL training, named as pill construction, pill poisoning, and pill injection accordingly. Extensive experimental results show that FL poisoning attacks enhanced by our method can bypass all the popular defenses, and can gain an up to 7x error rate increase, as well as on average a more than 2x error rate increase on both IID and non-IID data, in both cross-silo and cross-device FL systems.
1 Introduction
With the prosperity of machine learning and cloud computing, FL [1, 2] has been widely recognized as an effective approach to training machine learning models via distributed data on a large number of scattered clients. Compared with traditional centralized machine learning that collects data to a central cluster, FL does not require direct access to clients’ data, saving the communication overhead while protecting data privacy. However, due to its distributed architecture, FL is inherently vulnerable when clients are compromised by attackers. Numerous studies [3, 4, 5, 6, 7, 8] have demonstrated techniques that manipulate the behavior of trained global model via malicious clients, commonly referred to as poisoning attacks. These attacks can be further divided into the following two types: 1) Model poisoning attacks involve direct modifications of local model updates by the attacker [4, 6]. The corrupted updates, when aggregated, significantly skew the global model’s parameters to the opposite direction. 2) Data poisoning attacks inject malicious samples into the local training datasets [8, 9, 10, 11, 12, 13, 14, 15]. The compromised data samples lead to corrupted local model updates that, once merged, secretly deteriorate the global model’s integrity. Poisoning attacks expose severe security risks to FL [16, 17, 18, 19], violating its integrity and reliability.
To alleviate and further eliminate the impacts caused by poisoning attacks, researchers have proposed various defense methods. They can be divided into four main categories, including adaptive client filtering, statistical parameter aggregation, client-dominant detection, and other advanced metrics and pipelines. Adaptive client filtering [20, 21, 22, 23, 24, 25] utilizes various metrics or scores to evaluate the risk of clients, and discards updates from those with a high risk level. Statistical parameter aggregation [26, 27, 28, 29] modifies the standard aggregation process of FL, granting different weights to different clients based on their model update’s statistical features. Client-dominant detection [30, 31, 32, 33, 34] as a rising category leverages clients to maintain the integrity of FL training. While other advanced metrics and pipelines [35, 36, 37, 38, 39] use other advanced features or pipelines to perform the detection. The effectiveness of existing defense methods hinges on their ability to detect abnormal parameter updates, which are obvious when using most existing attacks. In particular, these attacks treat all neural network parameters equally and manipulate the entire model update using the same mutation strategy.
We argue that it is inefficient and ineffective for attackers to modify all the parameters simultaneously. Researches on model pruning [40, 41, 42, 43, 44] show that model parameters do not contribute equally to a model’s performance. Parameters that have little impact on a model’s performance are considered redundant. Manipulating these redundant parameters not only wastes the attacker’s budget but also decreases the stealthiness of the attack, especially when the FL server analyzes the overall statistics of model updates for anomaly detection. Thus, a more effective strategy is to manipulate critical parameters [45], which are crucial for the model’s performance. This approach ensures that the changes have a significant impact, enhancing the attack’s effectiveness while maintaining stealthiness.
We propose a novel attack-agnostic augmentation method in this paper, enhancing existing model poisoning attacks via a three-stage pipeline, including pill construction, pill poisoning, and pill injection. In the first stage – pill construction, we carefully design an abstract architecture of pill (a tiny subnet with a unique structure), named the pill blueprint, and identify the corresponding pill subnet instance in the target model via a dynamic search algorithm. Then in the second stage – pill poisoning, we reuse existing FL poisoning attacks in an attack-agnostic fashion and concentrate the malicious model updates to the selected pill subnet. Finally, in the pill injection stage, we insert the poisoned pill into a benign model update and use a two-step adjustment to further reduce the difference between the poisoned model update and the original benign update. To this end, we augment existing FL poisoning attacks by dynamically generating a pill, poisoning it, and gradually inserting it into the global model.
We conduct extensive experiments to evaluate the effectiveness of our augmentation method. We use our method to enhance four baseline poisoning attacks, namely, sign-flip attack, trim attack [4], krum attack [4], and min-max attack [6]. Using both the original and augmented attacks, we test the error rates (i.e., the proportion of incorrect predictions) of the global model trained by nine aggregation rules, including FedAvg [2], FLTrust [21], Multi-Krum [20], Median [26], Trim [26], Bulyan [27], FLDetector [39], DnC [6], and Flame [23]. The set of aggregation rules covers most existing defense metrics, to the best of our knowledge. We also design a corresponding adaptive defense to further test the robustness of our method when the defender has the full knowledge of our pipeline and detailed implementation. The experimental results illustrate that our method can substantially improve existing FL poisoning attacks by inducing a more than 2x increase on average in the model prediction error rate under existing defenses, and at most, a more than 7x error rate increase.
Our contributions can be summarized as follows:
  • We propose a generic and attack-agnostic augmentation approach designed to enhance poisoning attacks against robust FL. We are the first to encapsulate model poisoning attacks to well-defined subnets (i.e., pills) combined with comprehensive metric-based performance-boosting adjustment.
  • Extensive results against nine aggregation rules on three common datasets show the attack enhancement capability of our method. In particular, it helps four baseline model poisoning attacks bypass almost all the prevailing defense methods.
  • We observe and point out the inherent limitations of existing model poisoning attacks and defenses, exposing the necessity and potential of fine-grained FL security.
2 Background and Related Work
2.1 Federated Learning
Federated Learning (FL) [1, 2] trains a global model using the information from a swarm of clients without the direct access to each client’s data. In a standard FL training process, within an arbitrary communication round $t$, the FL server first distributes its global model $\boldsymbol{g}_{t}$ to all the clients $K$. After receiving this global model, each client $i$ trains a local model $\boldsymbol{g}_{t}^{(i)}$ with its local data $D^{(i)}$, and uploads the model update $\Delta\boldsymbol{g}_{t}^{(i)}$ to the FL server. After receiving the model updates from the clients, the FL server uses aggregation rules to calculate the global model $\boldsymbol{g}_{t+1}$ for the next round. The objective of FL can be formulated as:
$\mathop{min}_{\boldsymbol{g}}\sum_{i=0}^{K}\frac{|D^{(i)}|}{|D|}\cdot f(D^{(i)},\boldsymbol{g}).$ (1)
2.2 Poisoning Attacks in FL
Based on prior investigations [46, 47, 48], existing poisoning attacks in Federated Learning (FL) can be classified according to the techniques employed by attackers. Attackers may directly compromise the global model by manipulating the updates from local models [3, 4, 6, 7, 5] by compromised clients. Alternatively, they may poison their local datasets to indirectly influence the global model [9, 8, 10, 11, 12, 49]. The former technique is referred to as model poisoning attacks in FL, while the latter data poisoning attacks. Although model poisoning attacks are effective, existing attacks have limited stealthiness and can be detected by many existing defenses. Our goal is hence to demonstrate that such attacks can be augmented in a uniform way. Model poisoning attacks directly manipulate the parameters uploaded by clients, with a minimal interference to the local training process. Among these attacks, the simplest form is the sign-flipping attack, which directly flips the model update and scales it with a constant factor. A-Little-is-Enough [3] generates malicious updates within a calculated perturbation range to deceive the global model. Adaptive attacks [4], such as the Trim attack and the Krum attack, dynamically scale malicious updates based on parameter values and distances. The Min-Max and Min-Sum Attacks [6] provide a dynamic scaling for malicious updates based on different distance-based criteria. MPAF [7] aims to drive the global model towards a predefined target model with poor performance on given FL tasks. For the sake of generality, we employ the sign-flipping attack, two types of adaptive attacks [4] (the Trim attack and the Krum attack), and the Min-Max attack [6] as the baseline attacks in this paper.
Additionally, our method’s pill design is inspired by a specialized data poisoning attack known as the subnet replacement attack (SRA) [15]. This approach concentrates backdoor attacks within an arrow-width subnetwork of the original model. It trains this selected subnet using poisoned data and replaces the corresponding parameters of the target model with those from the trained subnetwork. Once the replacement is complete, SRA severs the connections between the poisoned subnetwork and the original model to preserve the efficacy of the attack. The stealthy yet effective design of SRA inspire our method. In particular, we devise a new subnet structure, referred to as the pill blueprint, which features heterogeneous widths to better accommodate a variety of existing FL poisoning attacks. Besides, unlike SRA’s one-time injection, our method gradually poisons the global model throughout the entire FL training, achieving better effectiveness against a wide range of defenses in the FL setting.
Figure 1: Overview of our augmentation method. The red parts indicate our augmentation method’s contribution, and the cyan parts represent the standard federated learning architecture.
2.3 Defenses against Poisoning Attacks in FL
Existing defenses can be categorized based on the mitigation strategies that they utilize.
Adaptive Client Filtering. These techniques such as Krum and Multi-Krum [20] filter out malicious clients through single or multiple rounds of client selection based on distance scores. FLTrust [21] computes trust scores using the cosine similarity between each client update and the server model’s update for weighted averaging. SignGuard [22] employs sign-based clustering combined with norm-based thresholding to identify and filter malicious clients. Flame [23] and Deepsight [24] propose adaptive clustering and clipping to safeguard against backdoor attacks. SkyMask [25] clusters trainable feature masks of clients to assess each client’s risk level.
Statistical Parameter Aggregation. Approaches like Median and Trim [26] use coordinate-wise median or trimmed mean values to aggregate model updates. Bulyan [27] enhances robustness by integrating Krum with Trim techniques. Fool’s Gold [28] applies an adaptive learning rate based on inter-client contribution similarity to mitigate the effects of malicious updates. SparseFed [29] aggregates sparsified updates, reducing the risk of model poisoning attacks.
Client-dominant Detection. Siren [30] and Siren+ [31] set proactive accuracy-based alarms at the client level with the corresponding server-side decisions to counter various model poisoning attacks. FL-WBC [32] introduces client-side noise to diminish the efficacy of attacks and shorten their duration. FLIP [33] achieves higher robustness through client-side reverse-engineering defenses against extensive poisoning strategies. LeadFL [34] uses a client-side Hessian matrix optimization to reduce the impact of adversarial patterns on backdoor and targeted attacks.
Other Advanced Metrics and Pipelines. Various studies employ other sophisticated metric pipelines designed for detection to ensure robust defense against poisoning attacks. These include techniques proposed in studies such as Zeno [35], CRFL [36], FedRecover [37], FLCert [38], FLDetector [39], and MESAS [50].
To comprhensively evaluate our method, we use Multi-Krum (MKrum) [20], Trimmed Mean (Trim) [26], Coordinate-wise Median (Median) [26], Bulyan [27], FLTrust [21], FLDetector (FLD) [39], DnC [6], and Flame [23], a set of representative methods, as our baselines. More details are shown in Appendix A.
3 Threat Model
Attacker’s Goal and Capabilities. This paper focuses on improving the effectiveness of existing poisoning attacks in FL. Similar to previous work [4, 6], an attacker aims to raise the error rates of the global model on a specific class or multiple classes by sending poisoned model updates via compromised clients during the iterative aggregation. Our method does not require any additional knowledge compared with existing FL poisoning attack. Hence we reuse the typical threat model in existing studies [4, 6]. The attacker has a complete control of the compromised clients, including their local data, local training, and uploading process. With the aggregated resources on the compromised clients, the attacker may aggregate the local data from the compromised clients to do extra training or aggregate their local updates to do model estimation. The attacker may or may not know the updates of other benign clients, depending on the confidentiality of the communication channels between the server and clients. Besides, the attacker cannot access the server’s information, including the aggregation rules or the selected clients in each round.
Defense Settings. Most of the defenses in FL are deployed and executed on the server. We adopt a similar defense setting as existing studies [20, 26, 21, 30]. The server cannot directly analyze the local data or the local training of clients. It can only detect malicious clients through model updates from different clients. The server can collect and possess a root test dataset to provide more accurate and robust detection, while the data of such a root test dataset cannot be derived from clients. The data distribution of this root test dataset may or may not be the same as the data distribution across the clients.
4 Design Objectives and Challenges
After analyzing the drawbacks and various implementations of existing FL poisoning attacks, we define three main objectives for our attack augmentation method: 1) For stealthiness, the augmentation method should stay stealthy while achieving comparable performance with original attacks. 2) For compatibility, the augmentation should be compatible with most of the existing FL poisoning attacks with few modifications on their implementations. 3) For generality, the attack augmentation should be able to bypass general detection methods with different detection metrics.
Corresponding to each objective, three challenges need to be addressed:
  • It presents a significant challenge that the attack augmentation method must use significantly fewer parameters while still achieving similar results as the original attacks.
  • It is challenging to develop a uniform augmentation method for various FL poisoning attacks since they require different information and are implemented in different training stages.
  • It is difficult to devise a general strategy that bypasses all common detection approaches, while guaranteeing the attack effectiveness.
5 Design
TABLE I: Main notations. Symbols in the gray part are used for attacks.
Symbol Meaning
$T$ Total FL communication round
$t$ FL communication round index
$K$ Total client number
$\boldsymbol{g}$ Global model of the FL training
$\boldsymbol{g}_{t}$ Global model in round $t$
$lr$ Learning rate
$f()$ Loss function used in the FL training
$i$ Client index
$E$ Local training epoch number
$D^{(i)}$ Local training data on client $i$
$\Delta\boldsymbol{g}_{t}^{(i)}$ Local model update of client $i$ in round $t$
$m$ Total amount of malicious clients
$D^{m}$ Aggregated data from compromised clients
$\Delta\hat{\boldsymbol{g}}_{t}^{m}$ Update of extra-trained model in round $t$
$\Delta\boldsymbol{\widetilde{g}}_{t}$ Estimated global model update in round $t$
$\Delta\boldsymbol{g}_{t}^{zero}$ Disconnection update in round $t$
$M$ Selected malicious subnetwork
$M_{disc}$ Disconnection mask corresponding to $M$
$C_{iter}$ Max malicious update adjustment iteration
$C_{\uparrow}$ Up-scaling factor
$C_{\downarrow}$ Down-scaling factor
1 function MalUpdate($i$,$t$,$\boldsymbol{g}_{t}$)
       % 1. Pill Construction
2       $\boldsymbol{M}$, $\boldsymbol{M}_{disc}\leftarrow$ Search($\boldsymbol{g}_{t}$);
       % 2. Pill Poisoning
3       $\boldsymbol{\hat{g}}^{m}_{t+1}\leftarrow\boldsymbol{g}_{t}$;
4       for each epoch $e_{extra}\leftarrow 1,\cdots,E_{extra}$ do
5             sample $B^{m}$ from aggregated local data $D^{m}$ on compromised clients ;
6             $\boldsymbol{\hat{g}}^{m}_{t+1}\leftarrow\boldsymbol{\hat{g}}^{m}_{t+1}-lr\cdot\nabla f(B^{m},\boldsymbol{\hat{g}}^{m}_{t+1}$);
7
8      $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}\leftarrow\boldsymbol{g}_{t}-\boldsymbol{\hat{g}}^{m}_{t+1}$;
9       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$ Poisoning($i$,$t$,$\boldsymbol{param}$,$\Delta\boldsymbol{\hat{g}}^{m}_{t+1}$);
10       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
       % 3. Poison Pill Injection
11       $\mathop{param}\leftarrow\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\}$;
12       $\Delta\boldsymbol{\widetilde{g}}_{t+1}\leftarrow$ Estimation($i$,$t$,$\boldsymbol{g}_{t}$,$\mathop{param}$);
13       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow\Delta\boldsymbol{g}_{t+1}^{(i)}+(\mathbf{1}-\boldsymbol{M})\odot\boldsymbol{\Delta\widetilde{g}}_{t+1}$;
14       $\Delta\boldsymbol{g}_{t+1}^{zero}\leftarrow\mathbf{0}-\boldsymbol{g}_{t}$;
15       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow\boldsymbol{M}_{disc}\odot\Delta\boldsymbol{g}_{t+1}^{zero}+(\mathbf{1}-\boldsymbol{M}_{disc})\odot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
16       $\mathop{param}=\mathop{param}\bigcup\{\boldsymbol{M}_{all}=M+M_{disc}\}$;
17       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
18       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
19       return $\Delta\boldsymbol{g}^{(i)}_{t+1}$
Algorithm 1 Our augmentation method’s workflow. Without losing generality, we assume the first $m$ clients as malicious clients.
5.1 Overview of Our Method
Fig. 1 presents the three key stages. Table I presents notations of main symbols in this paper.
[…]
We are the first that propose a universal attack augmentation pipeline for most FL poisoning attacks, considering stealthiness, compatibility, and generality. The detailed workflow of our method is shown in Algorithm 1.
5.2 Pill Construction
Figure 2: An example of the “approximate max pill search” algorithm in our augmentation method.
This stage aims to construct a pill structure for augmenting the stealthiness while retaining the attack effectiveness before being augmented. The pill is carefully crafted to involve a minimal subset of parameters from specific positions of the target model. We first define a pill’s blueprint as the pill’s graphic structure, independent of target model parameters. Then, we propose a dynamic pill search algorithm to identify and map concrete parameters from the target model to the blueprint.
Designing Pill Blueprint. The blueprint design is inspired by SRA [15], which shows that poisoning a narrow subnetwork (one neuron/channel in each layer) is adequate to effectively inject backdoors into machine learning models (not in the FL setting). However, their technique cannot be used for our purposes as their subnet architecture is very specific. It does not support attacking various targets; it is a fixed and pre-selected subnet without considering the dynamics of model training in FL; and its poisoned subnet is not stealthy, having substantially larger weight values compared to others due to the need to disseminate the poison effect through such a small pre-selected network. Therefore, we propose a novel blueprint method, in which the subnet structure is general, and its instantiations (i.e., the concrete subnets) vary across steps in the FL training procedure, including important neurons by a dynamic search algorithm. This allows small weight changes because poisoning important neurons enables easy dissemination, maximizing attack stealthiness. In particular, the pill blueprint is designed to accommodate various target classes of different FL poisoning attacks. It achieves this by manipulating the outputs relevant to multiple classes simultaneously, via disrupting all the output neurons together. Hence, our pill blueprint design follows the rules below:
  1. 1.
    The pill blueprint only contains one neuron in each linear layer or one channel in each convolutional layer, except for the last two layers. Suppose $\mathcal{N}_{i}^{p}$ represents the neuron/channel number in Layer $i$ in our pill blueprint, then $\mathcal{N}_{i}^{p}=1$ when $i<L-1$, where $L$ is the total layer counts in our pill blueprint.
  2. 2.
    In the last two layers of our pill blueprint, $\mathcal{N}^{p}_{L-1}=\mathcal{N}^{p}_{L}=$ number of classes.
Dynamic Pill Search. According to existing studies on neural network pruning [40, 41, 42, 43, 44], parameters with a larger magnitude typically dominate the model’s performance. The optimal solution is hence to examine the model parameters to search for a globally optimal pill that encompasses the most important parameters.
However, such a globally optimal pill could be identified via a pruning-based method [51, 52], and hence our attack could be easily detected. Besides, searching for a globally optimal pill is inefficient when the model has a large number of parameters. Thus, we search for an approximate pill instead, with an attacker-defined start point, and only evaluate a small subset of the entire model’s parameters. We name the search algorithm as “approximate max pill search”. A complete procedure of this search algorithm consists of the following four steps. To improve readability, we use “neuron” to represent both neurons in fully connected layers and channels in convolutional layers, and we use the classification task as an example:
Step 1 Random Start Point Selection:
At the beginning of the search, we randomly choose a subset of neurons from the first layer $l_{1}$ of the target model, based on the structure of the first layer $l^{p}_{1}$ in the pill’s blueprint and its neuron number $\mathcal{N}^{p}_{1}$. The selected neurons are termed as $\mathcal{V}_{1}$, and defined as start points, which are then fixed across the entire FL training.
Step 2 Layer-wise Search:
For each subsequent layer $l_{i}$ in the target model, we first calculate the sum of the weights from the selected neurons $\mathcal{V}_{i-1}$ in layer $l_{i-1}$ to each neuron in $l_{i}$. Then, we rank all the neurons in $l_{i}$ based on the parameter value sums and choose top $\mathcal{N}^{p}_{i}$ neurons in $l_{i}$ as the new $\mathcal{V}_{i}$, where $\mathcal{N}^{p}_{i}$ represents the number of neurons in the $i$th layer of pill’s blueprint. $\mathcal{V}_{i}$ and all the parameters from $\mathcal{V}_{i-1}$ to $\mathcal{V}_{i}$ are recorded.
Step 3 Output Neuron Pairing:
After visiting $l_{L-1}$ layers, where $L$ is the total number of layers in the target model, $||\mathcal{V}_{L-1}||$ should equal to the neuron number in $l_{L}$ of the target model, which also equals to the number of classes. We select all the neurons in the target model’s $l_{L}$ layer into our pill to construct $\mathcal{V}_{L}$. Then, we only record the parameters from one neuron in $\mathcal{V}_{L-1}$ to only one neuron in $\mathcal{V}_{L}$ based on the index order (i.e., the first neuron in $\mathcal{V}_{L-1}$ is paired with the first neuron in $\mathcal{V}_{L}$). Since $||\mathcal{V}_{L-1}||$ equals to $||\mathcal{V}_{L}||$, the number of recorded parameters equals to the number of classes, avoiding poisoning too many parameters in a single layer.
Step 4 Pill Mask Construction:
With the recorded $\mathcal{V}_{i}$ and the corresponding parameters, we construct two masks $\boldsymbol{M}$ and $\boldsymbol{M}_{disc}$. The mask $\boldsymbol{M}$ records the pill’s parameters in the target model, and the disconnection mask $\boldsymbol{M}_{disc}$ records the parameters of the connections between the pill and the rest of the target model. $\boldsymbol{M}$ is used for poisoning, while $\boldsymbol{M}_{disc}$ is used to disconnect the poison pill from the target model, maintaining the integrity and performance of the pill during poisoning. The two masks have the same shape as the target model’s parameters. To construct $\boldsymbol{M}$, we set the locations corresponding to the pill parameters to $1$, and the others to 00. Based on $\boldsymbol{M}$, we can similarly obtain the corresponding disconnected mask $\boldsymbol{M}_{disc}$, which sets the locations corresponding to the parameters from neurons other than $\mathcal{V}_{i-1}$ to $\mathcal{V}_{i}$ in each model layer $l_{i}$ (except for the $L$th layer since we choose all the neurons from it), and also those corresponding to parameters from $\mathcal{V}_{i-1}$ to other pill irrelevant neurons in each model layer $l_{i}$ to $1$. The two masks are used in the Pill Injection Stage.
Example. Fig. 2 presents a concrete example of the search algorithm in a 4-layer linear model. Since the start point is randomly selected by the attacker, defense methods can hardly guess it without any prior knowledge. In the example, suppose the model is a 4-layer linear models for a binary classification task. Then the pill blue print contains one neuron in the first two layers, and contains two neurons in the last two layers. Initially, we randomly select a start neuron, specifically the second neuron in the first layer in the example. Then, we conduct layer-wise searching when visiting the second and third layer, selecting the parameters with the highest magnitudes. At the forth layer (output layer), we pair the two output neurons with the two selected neurons in the third layer based on the index order. And we finally construct two pill-related masks accordingly.
With the search algorithm, we also reduce the complexity of the pill search from $\mathcal{O}(\prod_{1}^{L}\mathcal{N}_{i})$ to $\mathcal{O}(\sum_{1}^{L}\mathcal{N}_{i}\cdot\mathcal{N}^{p}_{i})$, where $\mathcal{N}_{i}$ represents the neuron number of the target model in layer $i$, and $\mathcal{N}^{p}_{i}\ll\mathcal{N}_{i}$ for all the hidden layers. The computational complexity of our pill search is hence much smaller than the computational complexity of one round local training.
TABLE II: Different dynamic patterns in our augmentation method, utilizing along with the max subnetwork searching.
Notation Description
$\textsc{Pattern}_{1}$ All layers use adaptive searching strategy
$\textsc{Pattern}_{2}$ All layers use one-time searching strategy
$\textsc{Pattern}_{3}$Centerstack[l]FE uses with adaptive searching strategy
CLS uses repeated searching strategy
$\textsc{Pattern}_{4}$Centerstack[l]FE uses repeated searching strategy
CLS uses with adaptive searching strategy
$\textsc{Pattern}_{5}$Centerstack[l]FE uses with adaptive searching strategy
CLS uses with one-time searching strategy
$\textsc{Pattern}_{6}$Centerstack[l]FE uses with one-time searching strategy
CLS uses with adaptive searching strategy
To make the pill search process dynamic, we also design several patterns to adaptively determine whether to change the pill in the training period, shown in Table II (FE represents the convolutional layers, CLS represents the linear layers). For more details about each specific dynamic pattern, please refer to Appendix C. The combination of the “approximate max pill search” with different dynamic patterns construct the complete dynamic pill search, considering both the stealthiness and the efficiency.
Figure 3: Cosine similarities between FLTrust server’s model update and malicious model update when malicious clients use different extra training rounds.
5.3 Pill Poisoning
In the pill poisoning stage, we aim to condense the poison into the pill using existing attacks. To achieve compatibility, our method simply reuses existing FL poisoning attacks, without any intrusive modification to their original implementations. We only modify the input of existing FL poisoning attacks by replacing the base model update with the update from a model that has undergone extra training rounds, denoted as $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$. Additionally, we restrict changes to parameters within the pill. The output is a poisoned pill that will be used in the next pill injection stage.
The motivation to use an extra-trained model update as the reference model update is shown in Fig. 3. As shown in the figure, with the increasing number of extra training rounds on the malicious clients, the generated malicious model update becomes less opposite to the FLTrust [21] server’s model update. Thus, we adopt the extra training in our method and limit the extra training epoch number $E_{extra}$ to less than the number of malicious clients $m$ times the benign local training epoch number $E$, denoted as $E_{extra}\leq m\cdot E$. With this extra training epoch number limit, we do not violate the threat model since the attacker can utilize the data and computational resources of all the compromised clients.
5.4 Pill Injection
In the pill injection stage, we aim to inject the pill into the model and use a two-step adjustment method to further camouflage the pill. Thus, the entire injection stage could be divided into two parts - pill injection and camouflaging. After this stage, the poison pill is seamlessly integrated with the benign model update and uploaded to the FL server.
Pill Insertion & Disconnection. In this part, our goal is to insert the pill into the model, and minimize the impact of the benign model updates on our pill. We use an estimated global model update as the benign model update, which is estimated as the coordinate-wise mean values of all the normal model updates from the compromised clients. The estimation process in the Estimation() in Algorithm 1 Line 11 is hence presented as Equations (2),
$\Delta\boldsymbol{\widetilde{g}}_{t+1}$ $\leftarrow\mathop{mean}\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\},$ (2)
where $\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}$ is the normal updates from the compromised clients. By aggregating information from multiple malicious clients, the estimated benign global model update is more similar to the genuine one, providing more budget for our poison pill.
After obtaining the estimated global model update $\boldsymbol{\Delta\widetilde{g}}_{t+1}$, we directly replace the parameters corresponding to the pill parameters (which have been poisoned in the previous stage) via the pill’s mask $\boldsymbol{M}$. Then, we replace the parameters that connect the pill and the other estimated global model updates with the disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$, using the disconnection mask $\boldsymbol{M}_{disc}$. The disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$ is calculated as $0-\boldsymbol{g}_{t}$, and is bounded by the maximum and minimum values of the reference model update $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$. The disconnection update can gradually change the parameters of the connections between the pill and rest of the model to 00, and finally isolates the poison pill from the global model, guaranteeing the attacking effects of the poison pill.
1 function SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
2       $\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1},\boldsymbol{M}_{all}\}\leftarrow\mathop{param}$;
[…]
4       $\mathop{iter}\leftarrow 0$;
5       while Sim($\Delta\boldsymbol{\widetilde{g}}_{t+1},\Delta\boldsymbol{g}_{t+1}^{(i)}$)<$S_{max}$& $\mathop{iter}<C_{iter}$ do
6             if $\mathop{iter}\%2$ then
7                   $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow(C_{\uparrow}\cdot(\mathbf{1}-\boldsymbol{M}_{all})+\boldsymbol{M}_{all})\odot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
8                   $\mathop{iter}\leftarrow\mathop{iter}+1$;
9
10            else
11                   $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow((\mathbf{1}-\boldsymbol{M}_{all})+C_{\downarrow}\cdot\boldsymbol{M}_{all})\odot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
12                   $\mathop{iter}\leftarrow\mathop{iter}+1$;
13
14
15      return $\Delta\boldsymbol{g}_{t+1}^{(i)}$;
16
17
18function DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
19       $\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1},\boldsymbol{M}_{all}\}\leftarrow\mathop{param}$;
20       $\mathop{Dist}_{max}\leftarrow\mathop{max}\{||\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||;i\in\{1,\cdots,m\}\}$;
21       $\mathop{Dist}\leftarrow||\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$;
22       if $||C_{\downarrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||<||C_{\uparrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$ then
23             $C_{dist}\leftarrow C_{\downarrow}$;
24
25      else
26             $C_{dist}\leftarrow C_{\uparrow}$;
27
28      while $\mathop{Dist}\geq\mathop{Dist}_{max}$& $||C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||\leq\mathop{Dist}$ do
29             $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
30             $\mathop{Dist}\leftarrow||\Delta\boldsymbol{g}_{t+1}^{(i)}-||$;
31
32      return $\Delta\boldsymbol{g}_{t+1}^{(i)}$;
33
Algorithm 2 Similarity-based and distance-based adjustment functions in the Poison Pill Injection stage.
Figure 4: Intuition behind distance-based adjustment in our augmentation method.
Pill Adjustment. After the injection, we use a two-step adjustment to further adjust the pill, improving the generality against multiple detection metrics simultaneously. In this stage, we consider two prevailing detection metrics – distance and cosine similarity. To increase the cosine similarity between the poisoned model update and the benign model update in our method, we balance the magnitudes of both the poison pill’s parameters and the other benign parameters. Similarly, to minimize the distance discrepancy between the poisoned and benign model updates, we adjust the magnitude of the entire poisoned model update. Thus, we first use the similarity-based adjustment, then use the distance-based adjustment to balance the effectiveness and stealthiness poisoned model update. This two-step adjustment is particularly effective when combined with our method, which selectively poisons only a tiny subset of the model’s parameters. By altering just a few parameters, our method preserves a substantial number of benign parameters, which are crucial for making effective adjustments. As a result, the poisoned model update can bypass a wide range of defenses since they are typically designed based on the combination or variants of distance and cosine similarity metrics, and they usually do not anticipate such a focused and minimal interference in the model parameters. The details of the two seamless adjustments are as follows:
Similarity-based Adjustment.
As shown in Line 1--12 of Algorithm 2, we first compute the maximum cosine similarity $S_{max}$ between the normal model updates from the compromised clients and the estimated global model update in the current round. Then, we iteratively and alternately reduce the magnitude of the poison pill’s parameters with the down-scaling factor $C_{\downarrow}$, and increase the magnitude of the rest estimated global model update’s parameters with the up-scaling factor $C_{\uparrow}$, until the cosine similarity between the entire poisoned model update and the estimated global model update is greater than $S_{max}$ or the adjustment total iteration is greater than the threshold $C_{iter}$.
Distance-based Adjustment.
In the Distance-based Adjustment (Line 13--24 of Algorithm 2), we reuse the up-scaling factor $C_{\uparrow}$ and the down-scaling factor $C_{\downarrow}$ to adjust the magnitude of the entire poisoned model update. The intuition behind the Distance-based Adjustment is shown in Fig. 4. We first calculate the maximum distance between the normal model updates from the compromised clients and the estimated global model update in the current round. We use this maximum distance $Dist_{max}$ as the threshold in the distance-based adjustment. Then, we further determine the scaling factor that should be used by applying the two scaling factors $C_{\uparrow}$ and $C_{\downarrow}$ separately to the poisoned model update $\Delta\boldsymbol{g}_{t+1}^{(i)}$. The scaling factor that reduces the distance between $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is chosen as the initial scaling factor in the subsequent iterative scaling. We stop the scaling until the distance between the $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is smaller than $Dist_{max}$, or the scaling factor begins to increase such distance (reach the limit of the scaling).
6 Evaluation
This section assesses how our method enhances the effectiveness of existing FL poisoning attacks from multiple perspectives. We begin by evaluating the Augmentation Effectiveness of our method against four FL poisoning attacks, using nine prevailing defenses across three datasets, detailed in Section 6.2. Subsequently, we visualize the Stealthiness of our method under two prevailing detection metrics, as discussed in Section 6.3. Lastly, an Generality Analysis of our method is presented, which includes tests on various proportions of malicious clients, tests on both cross-silo and cross-device settings, and evaluates the impact of different pill search rules, outlined in Section 6.4. Our method significantly enhances the capabilities of existing FL poisoning attacks, successfully bypassing all $9$ baseline defenses in over 90% of cases, and increasing the error rates by up to seven times compared to the original attacks’ error rates. Moreover, it demonstrates robustness across varying data distributions, model architectures, proportions of malicious clients, and pill search rules.
6.1 Evaluation Settings
Attack, Defense, and Framework settings. In our experiments, we typically set the malicious proportion to $20\%$. We implement $9$ baseline aggregation rules, including FedAvg, FLTrust, Multi-Krum, Bulyan, Median, Trim, FLDetector, DnC, and Flame. We use our method to augment $4$ existing model poisoning attacks, including sign-flipping attack, Trim attack, Krum attack, and Min-Max attack. These attacks are chosen for their representativeness in illustrating the effectiveness of our method. We configure a 50-client FL system for both the MNIST and Fashion-MNIST datasets. For the CIFAR-10 dataset, a 30-client FL system is used. Our framework accommodates both cross-silo and cross-device settings. The entire framework is based on PyTorch [53].
Model, Dataset, and Hyper-Parameters. In our experiments, we employ a four-layer Convolutional Neural Network (CNN) and a simplified version of AlexNet [54]. The structures of the models and their corresponding pill blueprints are detailed in Appendix B. We evaluate our method on three widely-used datasets: MNIST [55], Fashion-MNIST [56], and CIFAR-10 [57]. We use the CNN model on MNIST and Fashion-MNIST datasets, and the AlexNet [54] on CIFAR-10 dataset. Each experiment is repeated five times to ensure reliability, with the mean and standard deviation (std) of the results reported.
TABLE III: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
$\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003
Sign-flipping Attack 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
$\pm$0.023 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.002 $\pm$0.001 $\pm$0.003 $\pm$0.020 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.006 $\pm$0.003 $\pm$0.003
0.667 0.115 0.764 0.379 0.523 0.314 0.646 0.543 0.122 0.754 0.430 0.522 0.311 0.688
+ Poison Pill $\pm$0.089 $\pm$0.004 $\pm$0.049 $\pm$0.104 $\pm$0.091 $\pm$0.018 $\pm$0.061 $\pm$0.150 $\pm$0.006 $\pm$0.129 $\pm$0.057 $\pm$0.038 $\pm$0.038 $\pm$0.067
Trim Attack 0.243 0.109 0.139 0.146 0.174 0.179 0.116 0.332 0.120 0.201 0.163 0.231 0.238 0.124
$\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.006 $\pm$0.006 $\pm$0.003 $\pm$0.001 $\pm$0.022 $\pm$0.005 $\pm$0.018 $\pm$0.004 $\pm$0.008 $\pm$0.009 $\pm$0.003
0.618 0.576 0.638 0.284 0.453 0.219 0.115 0.668 0.517 0.687 0.292 0.473 0.223 0.222
+ Poison Pill $\pm$0.071 $\pm$0.057 $\pm$0.041 $\pm$0.040 $\pm$0.091 $\pm$0.010 $\pm$0.003 $\pm$0.033 $\pm$0.038 $\pm$0.036 $\pm$0.047 $\pm$0.047 $\pm$0.016 $\pm$0.128
Krum Attack 0.116 0.109 0.189 0.201 0.172 0.137 0.786 0.128 0.116 0.235 0.276 0.217 0.160 0.947
$\pm$0.002 $\pm$0.003 $\pm$0.022 $\pm$0.009 $\pm$0.008 $\pm$0.003 $\pm$0.087 $\pm$0.004 $\pm$0.003 $\pm$0.059 $\pm$0.003 $\pm$0.005 $\pm$0.003 $\pm$0.030
0.735 0.155 0.715 0.422 0.578 0.310 0.637 0.716 0.151 0.737 0.468 0.730 0.334 0.690
+ Poison Pill $\pm$0.032 $\pm$0.032 $\pm$0.132 $\pm$0.046 $\pm$0.057 $\pm$0.009 $\pm$0.074 $\pm$0.104 $\pm$0.004 $\pm$0.078 $\pm$0.017 $\pm$0.168 $\pm$0.031 $\pm$0.079
Min-Max Attack 0.183 0.110 0.431 0.330 0.183 0.218 0.825 0.269 0.125 0.619 0.434 0.255 0.278 0.831
$\pm$0.008 $\pm$0.002 $\pm$0.029 $\pm$0.015 $\pm$0.009 $\pm$0.009 $\pm$0.052 $\pm$0.026 $\pm$0.015 $\pm$0.050 $\pm$0.080 $\pm$0.012 $\pm$0.007 $\pm$0.049
0.702 0.303 0.668 0.327 0.514 0.314 0.778 0.629 0.320 0.612 0.406 0.547 0.376 0.822
+ Poison Pill $\pm$0.114 $\pm$0.201 $\pm$0.116 $\pm$0.074 $\pm$0.053 $\pm$0.047 $\pm$0.063 $\pm$0.114 $\pm$0.115 $\pm$0.040 $\pm$0.065 $\pm$0.072 $\pm$0.119 $\pm$0.036
IID and Non-IID Data Settings. Our method was assessed under both IID and non-IID data distributions to understand its performance across data heterogeneity. For IID data setting, we uniformly split all the training data into $K$ shards, and distribute each shard to a random client. For non-IID data setting, we utilize the non-IID degree $p$ as defined in prior studies [4, 30]. A higher $p$ indicates greater data heterogeneity among the clients. Specifically, when $p=0.1$, the data configuration is essentially IID. We set $p=0.5$ to to intensify the non-IID condition, under which we create and allocate $K$ non-IID data shards to all the clients, simulating a more realistic and challenging FL environment. Given that FLTrust necessitates a root dataset at the server, we select this dataset first from the available training data. Subsequently, we distribute the remaining data among the clients according to the aforementioned IID and non-IID rules. This approach ensures that there is no overlap between server’s data and client’s data.
Configurations of Dynamic Patterns in Our Method. As outlined in Section 5.2, we design six dynamic patterns for the pill search. We systematically evaluate all six patterns and present the results of the most effective strategy.
Evaluation Metrics. We use error rates – defined as the proportion of incorrect predictions – to evaluate attack effectiveness. Given that the model poisoning attacks discussed are all untargeted, higher error rates indicate more effective attacks. To assess the stealthiness of our method in delivering malicious updates, we employ two metrics: 1) cosine similarity score: measures alignment with the server’s model update in FLTrust; 2) distance score: used in Multi-Krum to evaluate the closeness of poisoned updates to benign updates.
TABLE IV: Error rates under cross-silo setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
No Attack 0.106 0.104 0.103 0.108 0.127 0.107 0.116 0.111 0.119 0.113 0.113 0.140 0.114 0.123
$\pm$ 0.003 $\pm$0.003 $\pm$0.003 $\pm$0.004 $\pm$0.001 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.001 $\pm$0.002 $\pm$0.005 $\pm$0.002 $\pm$0.004
Sign-flipping Attack 0.964 0.109 0.108 0.110 0.130 0.108 0.117 0.909 0.119 0.114 0.119 0.144 0.120 0.125
$\pm$0.017 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.001 $\pm$0.005 $\pm$0.045 $\pm$0.002 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.002
0.320 0.116 0.162 0.151 0.323 0.148 0.699 0.269 0.120 0.239 0.164 0.364 0.168 0.242
+ Poison Pill $\pm$0.080 $\pm$0.007 $\pm$0.027 $\pm$0.010 $\pm$0.029 $\pm$0.007 $\pm$0.082 $\pm$0.174 $\pm$0.003 $\pm$0.101 $\pm$0.013 $\pm$0.031 $\pm$0.012 $\pm$0.198
Trim Attack 0.112 0.111 0.111 0.115 0.132 0.114 0.116 0.125 0.115 0.121 0.125 0.153 0.122 0.122
$\pm$0.002 $\pm$0.005 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.006 $\pm$0.001 $\pm$0.004 $\pm$0.005 $\pm$0.002 $\pm$0.002
0.508 0.139 0.334 0.126 0.284 0.127 0.120 0.528 0.148 0.455 0.143 0.287 0.146 0.136
+ Poison Pill $\pm$0.128 $\pm$0.012 $\pm$0.120 $\pm$0.006 $\pm$0.040 $\pm$0.004 $\pm$0.003 $\pm$0.051 $\pm$0.018 $\pm$0.151 $\pm$0.003 $\pm$0.023 $\pm$0.004 $\pm$0.012
Krum Attack 0.107 0.108 0.114 0.123 0.141 0.112 0.668 0.116 0.117 0.124 0.138 0.173 0.122 0.410
$\pm$0.004 $\pm$0.005 $\pm$0.001 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.134 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.001 $\pm$0.005 $\pm$0.003 $\pm$0.352
0.183 0.118 0.283 0.161 0.362 0.146 0.631 0.428 0.127 0.280 0.187 0.415 0.182 0.704
+ Poison Pill $\pm$0.039 $\pm$0.003 $\pm$0.210 $\pm$0.015 $\pm$0.027 $\pm$0.008 $\pm$0.089 $\pm$0.190 $\pm$0.005 $\pm$0.064 $\pm$0.012 $\pm$0.057 $\pm$0.009 $\pm$0.091
Min-Max Attack 0.117 0.108 0.118 0.135 0.142 0.128 0.111 0.124 0.119 0.142 0.166 0.162 0.145 0.136
$\pm$0.004 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.009 $\pm$0.008 $\pm$0.004 $\pm$0.003 $\pm$0.012 $\pm$0.003 $\pm$0.007 $\pm$0.004 $\pm$0.004 $\pm$0.002
0.439 0.129 0.361 0.136 0.343 0.150 0.715 0.521 0.136 0.339 0.153 0.368 0.184 0.335
+ Poison Pill $\pm$0.140 $\pm$0.009 $\pm$0.245 $\pm$0.015 $\pm$0.032 $\pm$0.006 $\pm$0.096 $\pm$0.073 $\pm$0.011 $\pm$0.202 $\pm$0.009 $\pm$0.048 $\pm$0.017 $\pm$0.185
Figure 5: Comparison of Multi-Krum distance score between benign updates and malicious updates when using original poisoning attacks with and without our method.
TABLE V: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on CIFAR-10 dataset.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame
No Attack 0.488 0.480 0.507 0.469 0.551 0.456 0.445 0.494 0.491 0.486 0.474 0.499 0.498 0.581 0.502 0.463 0.506 0.532
Sign-flipping Attack 0.898 0.479 0.580 0.539 0.621 0.461 0.468 0.497 0.509 0.905 0.514 0.511 0.622 0.658 0.573 0.502 0.603 0.533
+ Poison Pill 0.739 0.880 0.929 0.694 0.707 0.699 0.536 0.899 0.706 0.879 0.861 0.898 0.677 0.766 0.688 0.566 0.900 0.675
Trim Attack 0.482 0.509 0.489 0.536 0.623 0.514 0.456 0.459 0.501 0.571 0.493 0.608 0.595 0.653 0.549 0.481 0.482 0.506
+ Poison Pill 0.853 0.877 0.883 0.654 0.674 0.662 0.513 0.899 0.542 0.890 0.862 0.906 0.772 0.688 0.639 0.518 0.893 0.621
Krum Attack 0.473 0.541 0.471 0.568 0.540 0.510 0.455 0.802 0.500 0.485 0.506 0.497 0.522 0.647 0.519 0.481 0.899 0.501
+ Poison Pill 0.701 0.896 0.900 0.765 0.756 0.643 0.529 0.890 0.872 0.724 0.849 0.900 0.675 0.748 0.647 0.580 0.885 0.873
Min-Max Attack 0.450 0.504 0.469 0.507 0.579 0.465 0.514 0.525 0.525 0.478 0.502 0.493 0.568 0.636 0.603 0.478 0.482 0.488
+ Poison Pill 0.752 0.712 0.902 0.775 0.802 0.640 0.545 0.902 0.811 0.661 0.646 0.886 0.674 0.783 0.661 0.527 0.907 0.799
Figure 6: Comparison of cosine similarity scores between original attack with and without our method under FLTrust.
Figure 7: Comparison of the cosine similarity scores with the server model of the original attack and our different augmentation patterns in the entire training period under FLTrust.
Figure 8: Comparison of error rates between original poisoning attacks and our method with two different pill search methods.
6.2 Augmentation Effectiveness
In this section, we present a comprehensive analysis of our method’s augmentation effectiveness on Fashion-MNIST dataset within a 50-client cross-silo FL system, in which 20% clients are malicious. We evaluate our method on both IID data and non-IID data. Our method successfully augments all the baseline attacks with more than $0.25$ average error rate increase, showing our method’s effectiveness and high compatibility.
Results on IID Data. The error rates of four baseline FL poisoning attacks, with and without our method, are shown in the left half of Table III. Our method enhances the error rates of the existing poisoning attacks in $23$ out of $28$ scenarios, against FedAvg and five defenses. The maximum increase in error rate is $0.658$, and the average increase reaches $0.274$. This substantial elevation from the attack-free baseline error rate of $0.109$ underscores our method’s capability to significantly compromise existing defenses’ integrity. Following are individual improvements of our method on different baseline attacks:
  • Sign-flipping attack: Its original version achieves a high error rate due to its aggressive and brute design, but it is effective only under FedAvg. Our method extends its impact to five more defenses (Multi-Krum, Bulyan, Median, Trim, and FLD), raising the average error rate by $0.399$.
  • Trim and Krum attack: Our method enables these two attacks to successfully penetrate all baseline defenses (except for Trim attack against FLD) including FLTrust, which were previously unbreachable, with average error rate increases of $0.249$ and $0.253$, respectively.
  • Min-Max attack: With our method, the Min-Max Attack shows a comprehensive improvement against all defenses except for a slight decrease against Bulyan, achieving an overall average error rate increase of $0.222$.
Results on Non-IID Data. Evaluations on non-IID data further validate the effectiveness of our method, demonstrating its superiority in $23$ out of $28$ cases. The highest error rate increase reaches $0.637$, with an average increase of $0.281$. Although there is a slight reduction in maximal error rate increase in the non-IID data setting, these results still demonstrate our method’s ability to effectively enhance attacks in more complex and heterogeneous data environments. The detailed improvements for a specific attack are shown as follows:
  • Sign-flipping attack: Our method helps the sign-flipping attack achieve an average error rate increase of $0.404$, which is similar to the error rate increase on IID data.
  • Trim and Krum attack: Both attacks penetrate all baseline defenses under the enhancement of our method, with average improvements of $0.281$ and $0.236$, respectively.
  • Min-Max attack: Our method helps the Min-Max attack achieve an average error rate increase of $0.195$, higher than its original version. Although this error rate increase is lower than that in the IID data setting, it remains higher than the error rate increase caused by its original version.
All attacks augmented by our method can bypass all baseline defenses, including FLTrust and FLDetector, with the exception of the sign-flipping attack. Notably, the Min-Max attack demonstrates superior effectiveness in non-IID data settings, achieving significant improvements compared to its performance on IID data. Other attacks also exhibit similar error rate improvements relative to their results on IID data, indicating that our method maintains its robustness and effectiveness in more complex data environments.
6.3 Stealthiness Analysis
To further analyze the performance of our method, we analyze its stealthiness during the training process of the FL system, focusing on how our method influences the distance scores and cosine similarity scores of exisitng FL poisoning attacks. The results indicate that our method can make malicious clients appear as benign or even more “benign” than genuine benign clients. This significant increase in stealthiness is a result of the pill design with distance-based adjustment and similarity-based adjustment techniques in our method.
Distance Score Analysis. Figure 5 compares the average distance scores of benign and malicious clients (with and without our method) across four baseline model poisoning attacks. The distance scores when using our method closely match or are even identical to those of benign clients throughout the entire training process. In contrast, original attacks like the Trim and sign-flipping attacks display distance scores that were significantly higher or lower than those of benign updates, indicating either detected by Multi-Krum (higher scores) or underutilized attack capacities (lower scores). Our method also has lower distance score variance in the early FL training period, representing that our method provides more steady attack efficacy in the FL’s critical training period [58, 59] by fully utilizing the attack capacities while being undetected. More details are shown in Appendix D.
Cosine Similarity Score Analysis. Figure 6 shows that the angles between server model updates and malicious updates using our method are similar or even smaller than those of benign updates, leading to higher aggregation weights for malicious updates in FLTrust – illustrating why our method makes existing FL poisoning attacks effectively bypass FLTrust. In contrast, the angles between the FLTrust’s server model updates with original malicious updates are often greater than $90^{\circ}$, leading to a zero aggregation weight. Detailed per-round cosine similarity trends (Figure 7) also reveal that while original attacks often result in negative similarities (and thus are excluded by FLTrust), our method maintains positive similarities throughout the entire training process. This consistency not only ensures the successful insertion of the pill in any specific round but also secures pill’s long-lasting presence in the global model.
6.4 Generality Analysis
In this section, we further discuss the generality of our method from four perspectives: malicious client proportion, client participation frequency, datasets & model architectures, and pill search algorithm. The results indicate that our method maintains its augmentation effectiveness consistently, even as these conditions change, demonstrating its reliability and wide applicability in augmenting FL poisoning attacks.
Impact of The Malicious Client’s Proportion. We first assess the effectiveness of our method in both IID and non-IID cross-silo FL system with only $10$% of clients compromised, as shown in Table IV. This setup revealed that all baseline model poisoning attacks yield lower error rates on the global model compared with scenarios with $20$% compromised clients. While the increase in error rates was less than those in the $20$% compromised client scenario, our method still effectively raises the global model’s error rates in $25$/$26$ out of $28$ cases (IID/non-IID setting). The maximum increase in error rates reached $0.403$, with an average increase of $0.144$. This average is notably higher (>$2$x higher) than the error rates observed in attack-free FL conditions. Specifically, our method help sign-flipping/Trim/Krum/Min-Max attacks achieve an average error rate increase of $0.133$/$0.094$/$0.136$/$0.209$. More detailed results are presented in Appendix G.
Impact of The Client Participation Frequency. We then extend the evaluation of our method to a cross-device FL system, where only $40$% of clients are selected for participation in each communication round. This setup results in less frequent participation from each client and a fluctuating proportion of malicious clients across different rounds. The maximum error rate increase with our method is $0.639$, with an average increase across different attacks and defenses of $0.279$. These results are consistent with those from the cross-silo FL system, underscoring our method’s effectiveness and generality across different FL configurations. This evaluation demonstrates our method’s robust performance and adaptability, not only in a controlled cross-silo environment but also under the more various conditions in cross-device FL systems. More details are presented in Appendix F.
Impact of The Datasets and Model Architectures. Following the evaluation with the Fashion-MNIST dataset, we test our method on the MNIST and CIFAR-10 dataset, employing the four-layer CNN model and the AlexNet model to further verify our method’s generality across different datasets. The collective results show that our method performs even better with larger datasets or more complex machine learning models. This trend confirms the generality of our method by revealing its capability to maintain consistent performance enhancements regardless of the dataset or model complexity involved. Specifically, our method help all the four baseline attacks bypass all the nine baselines on CIFAR-10 dataset, achieving $0.288$ error rate increase on average, presented in Table V. More detailed results on MNIST dataset are shown in Appendix E.
Impact of The Pill Search Algorithm. We conduct a final evaluation to assess the importance and effectiveness of the "approximate max pill search" algorithm used in our method. This is contrasted against a newly devised "approximate min pill search" algorithm, which targets the least important parameters within the target model. Figure 8 illustrates the error rates achieved by the "approximate max pill search", the "approximate min pill search", and the original model poisoning attacks. The "approximate max pill search" algorithm outperforms the "approximate min pill search" in $41$ out of $56$ cases (approximately $73$%), underscoring its effectiveness in leveraging the most influential parameters to enhance attack impacts. Despite its lower efficacy, the "approximate min pill search" still manage to surpass the original attacks in $41$ out of $56$ cases (approximately $73$%). This demonstrates the generality of our method across different pill search algorithms.
7 Discussion
7.1 Our Method against Possible Adaptive Defense
To further evaluate the robustness of our method when defenses are aware of the attack strategies (white-box scenario), we develop an adaptive defense named DSTrust, which enhances the FLTrust’s mechanism. DSTrust incorporates both distance and cosine similarity scores into a unified trust score calculation, directly countering our method’s two-step adjustment approach. The round-$t$ trust score of client $i$ in DSTrust is calculated as follows:
$TS_{i}=ReLU(\frac{\mathop{cos}(\Delta\boldsymbol{g}_{t}^{(i)},\Delta\boldsymbol{g}_{t}^{s})}{||\Delta\boldsymbol{g}_{t}^{(i)}-\Delta\boldsymbol{g}_{t}^{s}||}),$ (3)
where $\Delta\boldsymbol{g}_{t}^{(i)}$ represents the model update from client $i$ and $\Delta\boldsymbol{g}_{t}^{s})$ represents server’s model update. By integrating both cosine similarity and distance metrics, DSTrust provides a more comprehensive defense approach compared with FLTrust. This dual consideration allows DSTrust to effectively mitigate attacks that manipulate either of these metrics to bypass defenses.
TABLE VI: Error rates under cross-silo setting against the new adaptive defense – DSTrust – with and without our method on Fashion-MNIST dataset (20% malicious clients).
Data Distribution IID Non-IID
[…]
 Poison Pill
No Attack 0.108 0.116
[…]
 Attack 0.111 0.129 0.110 0.131
Trim Attack 0.109 0.629 0.115 0.630
Krum Attack 0.111 0.140 0.120 0.128
Min-Max Attack 0.127 0.167 0.143 0.327
Table VI details the error rates for four baseline FL poisoning attacks both with and without our method against the DSTrust defense on the Fashion-MNIST dataset within a 50-client FL system, where 20% of clients are malicious. These tests were conducted under both IID and non-IID data environments. DSTrust effectively neutralizes the four baseline poisoning attacks when our method is not applied, highlighting its robustness as a defense mechanism. Despite DSTrust’s integration of both cosine similarity and distance metrics in its defense strategy, it fails to counteract the augmented attacks when our method is employed. Notably, our method achieves a maximum error rate increase of $0.521$, and an average error rate increase of $0.173$ across all $8$ test scenarios. These results demonstrate that merely understanding the adjustment strategies of our method, and subsequently integrating corresponding defense metrics, does not fundamentally negate the effectiveness of our method. Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics into DSTrust, it remains insufficient to thwart the enhanced capabilities of our method.
7.2 Limitations and Future Work
Our method significantly enhances non-state-of-the-art (non-SOTA) model poisoning attacks, enabling them to SOTA results against various prevalent defenses. This is accomplished through a pill-based, attack-agnostic augmentation pipeline. We not only demonstrate our method’s capabilities but also expose fundamental vulnerabilities within the current designs of defense mechanisms.
For future attacks in FL, it is essential for attackers to meticulously evaluate the importance of each parameter in their implementation. By targeting specific subsets of parameters, attackers can devise more flexible and adaptive attacks, improving stealthiness and complicating defense efforts. As for future defenses, while individually checking each parameter might seem viable, its practical deployment is hindered by high overheads, making it infeasible in real-world applications.
Thus, there is a pressing need for more sophisticated defenses that can conduct fine-grained analyses of the roles of different parameters in neural networks, while executing without imposing prohibitive computational costs.
8 Conclusion
In this paper, we propose a novel attack-agnostic augmentation method to enhance existing poisoning attacks in FL by concentrating attacks into a pill (a tiny subnet). Our approach is constructed with three stages, including pill construction, pill poisoning, and pill injection. Accordingly, we first use a dynamic pill search algorithm to identify the pill blueprint. Then we poison the pill using existing FL poisoning attacks, and carefully inject the poison pill into the target model with two pill-related masks and a two-step adjustment. Our method enables existing poisoning attacks to achieve more than $2$x the error rates on average compared with their original implementations. The effectiveness of our method in exploiting and exacerbating the inherent weaknesses of current FL defenses highlights the critical need for more refined detection measures for FL.
Appendix A Additional Details of The Baseline Defenses
Krum and Multi-Krum (MKrum) [20]. Krum uses a distance score as the metric. In each round, the Krum server sums the distances between each client update $\boldsymbol{g}_{t}^{(i)}$ and its $K-m-2$ neighbors, and uses these sums as the scores for all the clients. The Krum server then selects the client’s model update with the lowest score. Multi-Krum is a variant of Krum that uses iterative Krum to pick multiple candidates for aggregation.
Coordinate-wise Median (Median) [26]. Coordinate-wise Median (Median) uses the per-parameter median values of the model updates from the clients as the aggregated global model update, which is then used to generate the next-round global model.
Trimmed Mean (Trim) [26]. Trimmed Mean (Trim) calculates per-parameter trimmed mean values of the client model updates and packs them as the global model update.
Bulyan [27]. Bulyan is a combination of Krum and Trim. It first uses the Krum-based method to select multiple candidates, and uses the per-parameter trimmed mean values of the candidate model updates as the final global model update.
FLTrust [21]. FLTrust trains a server model with a small root dataset. In each round, it computes the clipped cosine similarities between the server model update and client updates as trust scores, and then uses the trust scores as weights to aggregate all the normalized client model updates.
FLDetector (FLD) [39]. FLDetector filters out malicious clients by checking the multi-round consistency of all client updates. Malicious updates typically have lower consistency compared to benign ones.
Flame [23]. Flame utilizes HDBSCAN-based[60] dynamic clustering to filtered out malicious clients, and aggregates median-clipped benign updates with adaptive noise as the global model update.
Appendix B Additional Details of Concrete Pill Blueprints
Table VII and Table VIII illustrate the model structures of the CNN model and the simplified AlexNet, with their corresponding pill’s blueprints.
TABLE VII: Architectures of the original CNN model and the corresponding pill blueprint.
[…]
Pill Blueprint
Input $28\times 28\times 1$ $28\times 28\times 1$
Conv2d $3\times 3\times 30$ $3\times 3\times 1$
ReLU -1 -
MaxPool2d $2\times 2$ $2\times 2$
Conv2d $3\times 3\times 50$ $3\times 3\times 1$
ReLU - -
MaxPool2d $2\times 2$ $2\times 2$
Linear $1250\times 100$ $25\times 10$
ReLU - -
Linaer $100\times 10$ $10$
[…]
  • 1
    “-” represents that the model has this layer with no specified configuration.
  • 2
    $\times$” represents that the model does not contain this layer.
TABLE VIII: Example architectures of original AlexNet and the corresponding pill blueprint.
[…]
Pill Blueprint
Input $32\times 32\times 3$ $32\times 32\times 3$
Conv2d $11\times 11\times 48$ $11\times 11\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Conv2d $3\times 3\times 96$ $3\times 3\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Conv2d $3\times 3\times 192$ $3\times 3\times 1$
ReLU - -
Conv2d $3\times 3\times 192$ $3\times 3\times 1$
ReLU - -
Conv2d $3\times 3\times 128$ $3\times 3\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Linear $4608\times 1024$ $36\times 1$
ReLU - -
Linear $1024\times 512$ $1\times 10$
ReLU - -
Linear $512\times 10$ $10$
Softmax - $\times$
Appendix C Additional Details of The Dynamic Patterns in Our Method
We first design three searching strategies, including one-time searching strategy, repeated searching strategy, and adaptive searching strategy.
In the one-time searching strategy, we search the pill based on the initial global model using the “approximate max pill search” algorithm introduced in §5.2, and keep this pill unchanged in the whole FL training. The one-time searching strategy benefits the formation of pill in the global model, while the initial pill may be less effective with the increasing training rounds of the global model, due to the changing importance of the model parameters.
On the contrary, the repeated searching strategy runs the ‘approximate max pill search” algorithm in every training round. The repeated searching strategy can help our method modify more parameters in the global model, and make the pill less traceable. While the attacking effects may be reduced due to the constantly changing pill.
Considering advantages and disadvantages of both the one-time searching strategy and the repeated searching strategy, we design a more flexible searching strategy, termed as “adaptive searching strategy”. In adaptive searching strategy, our method searches the new pill only when the pill is not successfully injected into the global model in the last round. The condition:
[…]
should be satisfied to trigger the new subnetwork searching on malicious client $i$, where $C_{search}$ is set as $0.94$ in the experiments. The adaptive searching strategy is a mroe moderate version of repeated searching.
Since the three searching strategies have their unique advantages, we investigate different combinations of them in the experiments. We further divide the neural network into Feature Extractor (FE) and Classifier (CLS). Refer to the CNN model we used, the convolutional layers are regarded as FE, and the linear layers are regarded as CLS. We use different searching strategies in FE and CLS, respectively. In all the nine combinations, we test and keep six of them, noted as Pattern1 to Pattern6, shown in Table II in § 5.2. Such six patterns construct the entire dynamic pattern set used in our method.
TABLE IX: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on MNIST dataset.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
No Attack 0.028 0.051 0.029 0.029 0.045 0.029 0.025 0.029 0.042 0.030 0.029 0.041 0.029 0.022
Sign-flipping Attack 0.934 0.059 0.038 0.055 0.055 0.036 0.025 0.886 0.073 0.041 0.052 0.059 0.041 0.026
+ Poison Pill 0.353 0.093 0.454 0.283 0.268 0.173 0.588 0.431 0.059 0.605 0.349 0.333 0.217 0.713
Trim Attack 0.257 0.065 0.182 0.103 0.106 0.123 0.022 0.418 0.059 0.295 0.209 0.245 0.310 0.021
+ Poison Pill 0.416 0.109 0.469 0.252 0.247 0.117 0.026 0.581 0.065 0.672 0.358 0.324 0.092 0.051
Krum Attack 0.033 0.061 0.067 0.154 0.188 0.043 0.759 0.034 0.058 0.130 0.297 0.191 0.052 0.908
+ Poison Pill 0.326 0.082 0.585 0.266 0.272 0.169 0.632 0.528 0.062 0.556 0.350 0.321 0.210 0.746
Min-Max Attack 0.307 0.082 0.693 0.731 0.341 0.255 0.915 0.359 0.161 0.718 0.993 0.381 0.320 0.853
+ Poison Pill 0.402 0.106 0.518 0.273 0.262 0.218 0.766 0.534 0.077 0.707 0.369 0.318 0.194 0.861
TABLE X: Error rates under cross-device setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
No Attack 0.107 0.111 0.108 0.105 0.138 0.106 0.113 0.124 0.115 0.118 0.164 0.116
$\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.008 $\pm$0.006 $\pm$0.002 $\pm$0.009 $\pm$0.005
Sign-flipping Attack 0.940 0.116 0.110 0.128 0.165 0.121 0.905 0.124 0.118 0.136 0.184 0.134
$\pm$0.026 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.007 $\pm$0.004 $\pm$0.031 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.007 $\pm$0.006
0.591 0.117 0.749 0.357 0.589 0.225 0.573 0.125 0.665 0.379 0.662 0.277
+ Poison Pill $\pm$0.177 $\pm$0.004 $\pm$0.076 $\pm$0.057 $\pm$0.048 $\pm$0.026 $\pm$0.140 $\pm$0.004 $\pm$0.111 $\pm$0.018 $\pm$0.131 $\pm$0.012
Trim Attack 0.240 0.110 0.151 0.148 0.207 0.178 0.340 0.120 0.228 0.190 0.237 0.245
$\pm$0.018 $\pm$0.004 $\pm$0.010 $\pm$0.002 $\pm$0.014 $\pm$0.004 $\pm$0.048 $\pm$0.002 $\pm$0.025 $\pm$0.016 $\pm$0.016 $\pm$0.011
0.620 0.492 0.620 0.228 0.424 0.232 0.654 0.533 0.679 0.324 0.483 0.226
+ Poison Pill $\pm$0.051 $\pm$0.023 $\pm$0.025 $\pm$0.025 $\pm$0.042 $\pm$0.035 $\pm$0.037 $\pm$0.041 $\pm$0.049 $\pm$0.098 $\pm$0.098 $\pm$0.025
Krum Attack 0.117 0.112 0.172 0.238 0.169 0.132 0.126 0.121 0.204 0.296 0.222 0.158
$\pm$0.002 $\pm$0.004 $\pm$0.010 $\pm$0.005 $\pm$0.009 $\pm$0.003 $\pm$0.005 $\pm$0.004 $\pm$0.031 $\pm$0.011 $\pm$0.014 $\pm$0.006
0.681 0.138 0.740 0.362 0.572 0.258 0.604 0.141 0.750 0.372 0.649 0.277
+ Poison Pill $\pm$0.057 $\pm$0.015 $\pm$0.092 $\pm$0.073 $\pm$0.167 $\pm$0.018 $\pm$0.125 $\pm$0.009 $\pm$0.081 $\pm$0.035 $\pm$0.184 $\pm$0.013
Min-Max Attack 0.146 0.111 0.382 0.324 0.183 0.185 0.191 0.147 0.621 0.426 0.245 0.279
$\pm$0.005 $\pm$0.002 $\pm$0.036 $\pm$0.012 $\pm$0.011 $\pm$0.006 $\pm$0.014 $\pm$0.024 $\pm$0.112 $\pm$0.072 $\pm$0.013 $\pm$0.007
0.651 0.244 0.718 0.312 0.503 0.249 0.670 0.229 0.621 0.349 0.581 0.386
+ Poison Pill $\pm$0.082 $\pm$0.104 $\pm$0.059 $\pm$0.026 $\pm$0.060 $\pm$0.014 $\pm$0.123 $\pm$0.098 $\pm$0.030 $\pm$0.047 $\pm$0.161 $\pm$0.141
TABLE XI: Error rates under cross-device setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
No Attack 0.110 0.106 0.107 0.108 0.139 0.106 0.115 0.117 0.115 0.117 0.164 0.112
$\pm$0.004 $\pm$0.004 $\pm$0.005 $\pm$0.003 $\pm$0.008 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.002
Sign-flipping Attack 0.929 0.111 0.108 0.111 0.153 0.117 0.902 0.118 0.115 0.120 0.175 0.134
$\pm$0.026 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.025 $\pm$0.004 $\pm$0.034 $\pm$0.001 $\pm$0.004 $\pm$0.003 $\pm$0.008 $\pm$0.008
0.195 0.114 0.170 0.138 0.347 0.137 0.330 0.124 0.165 0.148 0.483 0.161
+ Poison Pill $\pm$0.032 $\pm$0.003 $\pm$0.071 $\pm$0.005 $\pm$0.059 $\pm$0.004 $\pm$0.135 $\pm$0.006 $\pm$0.016 $\pm$0.007 $\pm$0.225 $\pm$0.009
Trim Attack 0.112 0.114 0.111 0.118 0.153 0.113 0.129 0.125 0.128 0.129 0.185 0.122
$\pm$0.003 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.021 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.011 $\pm$0.003
0.369 0.138 0.212 0.128 0.310 0.140 0.589 0.154 0.300 0.139 0.351 0.156
+ Poison Pill $\pm$0.147 $\pm$0.015 $\pm$0.066 $\pm$0.005 $\pm$0.038 $\pm$0.014 $\pm$0.046 $\pm$0.017 $\pm$0.100 $\pm$0.006 $\pm$0.056 $\pm$0.015
Krum Attack 0.110 0.113 0.115 0.128 0.144 0.113 0.121 0.116 0.123 0.135 0.183 0.120
$\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
0.164 0.117 0.157 0.143 0.371 0.142 0.229 0.126 0.249 0.146 0.374 0.157
+ Poison Pill $\pm$0.038 $\pm$0.003 $\pm$0.044 $\pm$0.010 $\pm$0.034 $\pm$0.005 $\pm$0.069 $\pm$0.002 $\pm$0.167 $\pm$0.004 $\pm$0.023 $\pm$0.005
Min-Max Attack 0.116 0.111 0.116 0.127 0.145 0.122 0.121 0.116 0.123 0.135 0.183 0.120
$\pm$0.002 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.006 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
0.351 0.124 0.299 0.135 0.343 0.146 0.342 0.138 0.292 0.148 0.417 0.166
+ Poison Pill $\pm$0.204 $\pm$0.019 $\pm$0.110 $\pm$0.004 $\pm$0.070 $\pm$0.019 $\pm$0.076 $\pm$0.018 $\pm$0.087 $\pm$0.009 $\pm$0.050 $\pm$0.010
Appendix D Additional Stealthiness Analysis
As shown in Figure 5, in addition to the minimal difference between the benign updates and the augmented malicious updates, our method also achieves two more improvements. First, our method causes the global model to degrade earlier compared to the original attacks, further demonstrating the effectiveness of our augmentation. Second, our method significantly increases the discrepancy between benign client updates as the communication rounds increase. While original attacks can bypass detection in some cases, the discrepancy between benign client updates remains steady, illustrating the lower impact of malicious clients. In contrast, our method consistently increases the discrepancy among benign clients, highlighting its penetrating effectiveness in its influence on benign clients’ local training.
Appendix E Additonal Details on MNIST and CIFAR-10 Dataset
The detailed results on MNIST and CIFAR-10 dataset are presented in Table IX (MNIST dataset) and Table V (CIFAR-10 dataset), respectively.
For MNIST dataset, the highest error rate increase achieved using our method is $0.518$, with an average increase of $0.121$. This average error rate increase is slightly lower compared with the improvement observed on the Fashion-MNIST dataset. Despite the reduced average error rate increase, it remains significant, especially considering the MNIST dataset’s lower baseline error rates (below 0.070).
On CIFAR-10 dataset, our method help existing FL poisoning attacks outperform their original versions in $71$ of the $72$ scenarios, with an average error rate increase of over $0.288$. Specifically, our method facilitates at least a $0.212$ increase in error rates against FLTrust, outperforming the results in the same settings on the Fashion-MNIST dataset.
Appendix F Additional Results in Cross-device FL System
After evaluating our method in the 50-client cross-silo FL system, we further test it in the 50-client cross-device FL system. Table X presents the error rates under the cross-device FL setting using the “approximate max pill search” algorithm on both IID and non-IID data. We report the highest error rates among the results of six dynamic patterns, with the malicious client proportion set to $20\%$. Since FLD is not typically designed for cross-device systems, we do not test it in this setting.
Results on IID Data. The highest error rate improvement with our method achieves $0.639$, and the average error rate increase with our method reaches $0.279$. With our method, existing model poisoning attacks outperform their original versions in $22$ out of the $24$ cases. The highest error rate improvement for the sign-flipping attack is $0.639$, with an average error rate increase of $0.279$. For the Trim attack and Krum attack, the highest error rate increases are $0.469$ and $0.568$, with average error rate increases of $0.264$ and $0.302$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.505$, with an average increase of $0.272$. These improvements are consistent with the error rates observed under the cross-silo FL setting using the "approximate max pill search" algorithm on IID data.
Results on non-IID Data. As for the results on non-IID data, the highest error rate improvement with our method achieves $0.546$, and the average error rate increase with our method reaches $0.273$. By using our method, existing model poisoning attacks outperform their original versions in 21 out of the 24 cases. The highest error rate improvement for the sign-flipping attack is $0.547$, with an average error rate increase of $0.282$. For the Trim attack and Krum attack, the highest error rate rises are $0.451$ and $0.546$, with average error rate rises of $0.312$ and $0.278$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.479$, with an average increase of $0.201$. These improvements are also aligned with the error rates observed under the cross-silo FL setting using the max subnetwork searching algorithm on non-IID data.
The average error rates of the global model in the cross-device FL system are lower than the error rates in the cross-silo FL system within $0.030$, illustrating our method’s generality over different data distribution and FL systems.
Appendix G Additional Results with Fewer Malicious Clients
We also test the error rate improvement of our method in both the IID and non-IID cross-device FL systems, with only $10\%$ malicious clients. The experimental results are shown in Table XI.
Results on IID Data with Fewer Malicious Clients. The highest error rate increment is $0.257$, with an average increment of $0.083$. The error rate increments in the cross-device FL system are smaller than those in the cross-silo FL system, as malicious clients may not be selected in every round. However, this reduction in improvement is acceptable since our method helps existing model poisoning attacks outperform their original versions in $23$ out of $24$ cases. Furthermore, when all existing attacks fail to bypass any defenses with $10\%$ malicious clients, our method enables the attacks to bypass all defenses. The superiority of our method is maintained even with $10\%$ compromised clients.
Results on Non-IID Data with Fewer Malicious Clients. The results on the non-IID data are similar to those on the IID data. The highest error rate increment is $0.460$, and the average error rate increment is $0.079$. Our method helps existing model poisoning attacks achieve higher error rates in $23$ out of $24$ cases, even in highly unstable and heterogeneous settings. These results demonstrate the generality and robustness of our method across different data distributions and client selection methods with only a small portion of malicious clients.

### v2

@@
 Abstract
-Without direct access to the client’s data, federated learning (FL) is well-known for its unique strength in data privacy protection among existing distributed machine learning techniques.
-However, its distributive and iterative nature makes FL inherently vulnerable to various poisoning attacks, including model poisoning attacks and data poisoning attacks.
-To counteract these threats, extensive defenses have been proposed to filter out malicious clients, using various detection metrics.
-Based on our analysis of existing attacks and defenses, we find that there is a lack of attention to model redundancy.
-In neural networks, various model parameters contribute differently to the model’s performance.
-However, existing attacks in FL manipulate all the model update parameters with the same strategy, making them easily detectable by common defenses.
-Meanwhile, the defenses also tend to analyze the overall statistical features of the entire model updates, leaving room for sophisticated attacks.
-Based on these observations, this paper proposes a generic and attack-agnostic augmentation approach designed to enhance the effectiveness and stealthiness of existing FL poisoning attacks against detection in FL, pointing out the inherent flaws of existing defenses and exposing the necessity of fine-grained FL security.
-Specifically, we employ a three-stage methodology that strategically constructs, generates, and injects poison (generated by existing attacks) into a pill (a tiny subnet with a novel structure) during the FL training, named as pill construction, pill poisoning, and pill injection accordingly.
-Extensive experimental results show that FL poisoning attacks enhanced by our method can bypass all the popular defenses, and can gain an up to 7x error rate increase, as well as on average a more than 2x error rate increase on both IID and non-IID data, in both cross-silo and cross-device FL systems.
+Federated learning (FL) protects data privacy by enabling distributed model training without direct access to client data.
+However, its distributed nature makes it vulnerable to model and data poisoning attacks.
+While numerous defenses filter malicious clients using statistical metrics, they overlook the role of model redundancy, where not all parameters contribute equally to the model and attack performance.
+Current attacks manipulate all model parameters uniformly, making them more detectable, while defenses focus on the overall statistics of client updates, leaving gaps for more sophisticated attacks.
+We propose an attack-agnostic augmentation method to enhance the stealthiness and effectiveness of existing poisoning attacks in FL, exposing flaws in current defenses and highlighting the need for fine-grained FL security.
+Our three-stage methodology, including pill construction, pill poisoning, and pill injection, injects poison into a compact subnet (i.e., pill) of the global model during the iterative FL training.
+Experimental results show that FL poisoning attacks enhanced by our method can bypass 8 state-of-the-art (SOTA) defenses, gaining an up to 7x error rate increase, as well as on average a more than 2x error rate increase on both IID and non-IID data, in both cross-silo and cross-device FL systems.
+Codehttps://github.com/MarkGHX/PoisonPill
 1 Introduction
-With the prosperity of machine learning and cloud computing, FL [1, 2] has been widely recognized as an effective approach to training machine learning models via distributed data on a large number of scattered clients.
-Compared with traditional centralized machine learning that collects data to a central cluster, FL does not require direct access to clients’ data, saving the communication overhead while protecting data privacy.
-However, due to its distributed architecture, FL is inherently vulnerable when clients are compromised by attackers.
-Numerous studies [3, 4, 5, 6, 7, 8] have demonstrated techniques that manipulate the behavior of trained global model via malicious clients, commonly referred to as poisoning attacks.
-These attacks can be further divided into the following two types: 1) Model poisoning attacks involve direct modifications of local model updates by the attacker [4, 6].
-The corrupted updates, when aggregated, significantly skew the global model’s parameters to the opposite direction.
-2) Data poisoning attacks inject malicious samples into the local training datasets [8, 9, 10, 11, 12, 13, 14, 15].
-The compromised data samples lead to corrupted local model updates that, once merged, secretly deteriorate the global model’s integrity.
-Poisoning attacks expose severe security risks to FL [16, 17, 18, 19], violating its integrity and reliability.
-To alleviate and further eliminate the impacts caused by poisoning attacks, researchers have proposed various defense methods.
-They can be divided into four main categories, including adaptive client filtering, statistical parameter aggregation, client-dominant detection, and other advanced metrics and pipelines.
-Adaptive client filtering [20, 21, 22, 23, 24, 25] utilizes various metrics or scores to evaluate the risk of clients, and discards updates from those with a high risk level.
-Statistical parameter aggregation [26, 27, 28, 29] modifies the standard aggregation process of FL, granting different weights to different clients based on their model update’s statistical features.
-Client-dominant detection [30, 31, 32, 33, 34] as a rising category leverages clients to maintain the integrity of FL training.
-While other advanced metrics and pipelines [35, 36, 37, 38, 39] use other advanced features or pipelines to perform the detection.
-The effectiveness of existing defense methods hinges on their ability to detect abnormal parameter updates, which are obvious when using most existing attacks.
-In particular, these attacks treat all neural network parameters equally and manipulate the entire model update using the same mutation strategy.
-We argue that it is inefficient and ineffective for attackers to modify all the parameters simultaneously.
-Researches on model pruning [40, 41, 42, 43, 44] show that model parameters do not contribute equally to a model’s performance.
-Parameters that have little impact on a model’s performance are considered redundant.
-Manipulating these redundant parameters not only wastes the attacker’s budget but also decreases the stealthiness of the attack, especially when the FL server analyzes the overall statistics of model updates for anomaly detection.
-Thus, a more effective strategy is to manipulate critical parameters [45], which are crucial for the model’s performance.
-This approach ensures that the changes have a significant impact, enhancing the attack’s effectiveness while maintaining stealthiness.
-We propose a novel attack-agnostic augmentation method in this paper, enhancing existing model poisoning attacks via a three-stage pipeline, including pill construction, pill poisoning, and pill injection.
-In the first stage – pill construction, we carefully design an abstract architecture of pill (a tiny subnet with a unique structure), named the pill blueprint, and identify the corresponding pill subnet instance in the target model via a dynamic search algorithm.
-Then in the second stage – pill poisoning, we reuse existing FL poisoning attacks in an attack-agnostic fashion and concentrate the malicious model updates to the selected pill subnet.
-Finally, in the pill injection stage, we insert the poisoned pill into a benign model update and use a two-step adjustment to further reduce the difference between the poisoned model update and the original benign update.
-To this end, we augment existing FL poisoning attacks by dynamically generating a pill, poisoning it, and gradually inserting it into the global model.
+With the rising demand for machine learning and cloud computing, Federated Learning (FL)(konevcny2016federated; mcmahan2017communication) has emerged as a key approach for training models on distributed data from scattered clients.
+Unlike centralized machine learning, FL avoids direct data access, reducing communication overhead and enhancing privacy.
+However, its distributed nature leaves it vulnerable when clients are compromised.
+Numerous studies (baruch2019little; fang2020local; bhagoji2019analyzing; shejwalkar2021manipulating; cao2022mpaf; bagdasaryan2020backdoor) have examined poisoning attacks, where malicious clients manipulate the global model.
+These fall into two categories: 1) Model poisoning which directly alters local updates to skew global parameters (fang2020local; shejwalkar2021manipulating); and 2) Data poisoning that injects malicious samples into local datasets (bagdasaryan2020backdoor; tolpegin2020data; xie2020dba; sun2019can; wang2020attack; chen2017targeted; liu2018trojaning; qi2022sra).
+Such attacks threaten FL’s integrity and reliability (lyu2020threats; kairouz2021advances).
+To mitigate these attacks, defenses have been proposed, including adaptive client filtering (krum; cao2020fltrust; xu2021signguard; nguyen2022flame; yan2023skymask), statistical parameter aggregation (yin2018trim; bulyan; foolsgold; panda2022sparsefed; han2023towards), client-dominant detection (guo2021siren; guo2024siren+; sun2021flwbc; zhang2023flip; zhu2023leadfl; park2023feddefender), and other advanced metrics and pipelines (xie2019zeno; xie2021crfl; cao2023fedrecover; cao2022flcert; zhang2022fldetector).
+These approaches aim to identify suspicious updates, which are usually evident in model poisoning attacks that uniformly alter parameters.
+We argue that modifying all parameters uniformly is not a cost-effective approach.
+Studies on model pruning (frankle2018lottery; lin2018deep; han2015deep; mugunthan2022fedltn; jiang2022model) show that parameters do not contribute equally to a model’s performance.
+Altering redundant parameters wastes resources and reduces attack stealthiness.
+A more effective strategy is to target critical parameters (zhang2023oblivion), which significantly impact performance, thereby increasing the attack’s effectiveness while maintaining stealthiness.
+Thus, we propose a novel attack-agnostic augmentation method that enhances model poisoning attacks using a three-stage pipeline: pill construction, pill poisoning, and pill injection.
+In the first stage, we design a pill blueprint and identify its corresponding subnet instance in the target model.
+During pill poisoning, existing FL attacks are applied in an attack-agnostic manner to poison the selected pill.
+Finally, in pill injection, the poisoned pill is inserted into an estimated benign update, and a two-step adjustment is used to minimize the difference between the poisoned and benign updates.
+This approach dynamically generates, poisons, and injects a pill into the global model, augmenting existing FL poisoning attacks.
 We conduct extensive experiments to evaluate the effectiveness of our augmentation method.
-We use our method to enhance four baseline poisoning attacks, namely, sign-flip attack, trim attack [4], krum attack [4], and min-max attack [6].
-Using both the original and augmented attacks, we test the error rates (i.e., the proportion of incorrect predictions) of the global model trained by nine aggregation rules, including FedAvg [2], FLTrust [21], Multi-Krum [20], Median [26], Trim [26], Bulyan [27], FLDetector [39], DnC [6], and Flame [23].
-The set of aggregation rules covers most existing defense metrics, to the best of our knowledge.
-We also design a corresponding adaptive defense to further test the robustness of our method when the defender has the full knowledge of our pipeline and detailed implementation.
-The experimental results illustrate that our method can substantially improve existing FL poisoning attacks by inducing a more than 2x increase on average in the model prediction error rate under existing defenses, and at most, a more than 7x error rate increase.
-Our contributions can be summarized as follows:
-  • We propose a generic and attack-agnostic augmentation approach designed to enhance poisoning attacks against robust FL.
-We are the first to encapsulate model poisoning attacks to well-defined subnets (i.e., pills) combined with comprehensive metric-based performance-boosting adjustment.
-  • Extensive results against nine aggregation rules on three common datasets show the attack enhancement capability of our method.
-In particular, it helps four baseline model poisoning attacks bypass almost all the prevailing defense methods.
-  • We observe and point out the inherent limitations of existing model poisoning attacks and defenses, exposing the necessity and potential of fine-grained FL security.
+We apply it to four baseline poisoning attacks: sign-flipping attack, trim attack (fang2020local), krum attack (fang2020local), and min-max attack (shejwalkar2021manipulating).
+Using both the original and augmented versions, we measure error rates (i.e., the proportion of incorrect predictions) of the global model trained with nine aggregation rules: FedAvg (mcmahan2017communication), FLTrust (cao2020fltrust), Multi-Krum (krum), Median (yin2018trim), Trim (yin2018trim), Bulyan (bulyan), FLDetector (zhang2022fldetector), DnC (shejwalkar2021manipulating), and Flame (nguyen2022flame).
+These aggregation rules represent most existing defense metrics.
+We also design an adaptive defense where the defender has full knowledge of our pipeline and implementation.
+Results show our method substantially improves existing FL poisoning attacks, leading to over a 2x average increase in model prediction error rates under existing defenses, and up to a 7x increase in some cases.
+Our contributions are summarized as follows:
+  • We propose a generic, attack-agnostic augmentation method that enhances poisoning attacks against robust FL by encapsulating model poisoning attacks into well-defined subnets (i.e., pills) with comprehensive metric-based adjustments.
+  • Extensive experiments on three common datasets against nine aggregation rules demonstrate that our method helps baseline attacks bypass almost all existing defenses, which cannot be attacked by original versions.
+  • We identify limitations of existing poisoning attacks and defenses in FL, highlighting the need and potential for fine-grained FL security.
 2 Background and Related Work
 2.1 Federated Learning
-Federated Learning (FL) [1, 2] trains a global model using the information from a swarm of clients without the direct access to each client’s data.
+Federated Learning (FL) (konevcny2016federated; mcmahan2017communication) trains a global model using the information from a swarm of clients without the direct access to each client’s data.
 In a standard FL training process, within an arbitrary communication round $t$, the FL server first distributes its global model $\boldsymbol{g}_{t}$ to all the clients $K$.
@@
 2.2 Poisoning Attacks in FL
-Based on prior investigations [46, 47, 48], existing poisoning attacks in Federated Learning (FL) can be classified according to the techniques employed by attackers.
-Attackers may directly compromise the global model by manipulating the updates from local models [3, 4, 6, 7, 5] by compromised clients.
-Alternatively, they may poison their local datasets to indirectly influence the global model [9, 8, 10, 11, 12, 49].
-The former technique is referred to as model poisoning attacks in FL, while the latter data poisoning attacks.
-Although model poisoning attacks are effective, existing attacks have limited stealthiness and can be detected by many existing defenses.
-Our goal is hence to demonstrate that such attacks can be augmented in a uniform way.
-Model poisoning attacks directly manipulate the parameters uploaded by clients, with a minimal interference to the local training process.
-Among these attacks, the simplest form is the sign-flipping attack, which directly flips the model update and scales it with a constant factor.
-A-Little-is-Enough [3] generates malicious updates within a calculated perturbation range to deceive the global model.
-Adaptive attacks [4], such as the Trim attack and the Krum attack, dynamically scale malicious updates based on parameter values and distances.
-The Min-Max and Min-Sum Attacks [6] provide a dynamic scaling for malicious updates based on different distance-based criteria.
-MPAF [7] aims to drive the global model towards a predefined target model with poor performance on given FL tasks.
-For the sake of generality, we employ the sign-flipping attack, two types of adaptive attacks [4] (the Trim attack and the Krum attack), and the Min-Max attack [6] as the baseline attacks in this paper.
-Additionally, our method’s pill design is inspired by a specialized data poisoning attack known as the subnet replacement attack (SRA) [15].
-This approach concentrates backdoor attacks within an arrow-width subnetwork of the original model.
-It trains this selected subnet using poisoned data and replaces the corresponding parameters of the target model with those from the trained subnetwork.
-Once the replacement is complete, SRA severs the connections between the poisoned subnetwork and the original model to preserve the efficacy of the attack.
-The stealthy yet effective design of SRA inspire our method.
-In particular, we devise a new subnet structure, referred to as the pill blueprint, which features heterogeneous widths to better accommodate a variety of existing FL poisoning attacks.
-Besides, unlike SRA’s one-time injection, our method gradually poisons the global model throughout the entire FL training, achieving better effectiveness against a wide range of defenses in the FL setting.
+Following prior studies (shejwalkar2022back; khan2023pitfalls; jere2020taxonomy), poisoning attacks in Federated Learning (FL) fall into two categories: model poisoning and data poisoning.
+In model poisoning attacks, adversaries compromise the global model by directly altering local model updates (baruch2019little; fang2020local; shejwalkar2021manipulating; cao2022mpaf; bhagoji2019analyzing).
+In data poisoning attacks, they corrupt local datasets to indirectly affect the global model (tolpegin2020data; bagdasaryan2020backdoor; xie2020dba; sun2019can; wang2020attack; zhang2022neurotoxin).
+Further details are in the extended version.
+Our pill design draws inspiration from the subnet replacement attack (SRA) (qi2022sra), a backdoor injection method that limits backdoors to a small subnetwork.
+SRA trains this subnet with tainted data, substitutes the target model’s parameters, and disconnects the subnet to maintain attack effect.
+Drawing from SRA’s stealthy yet potent design, we propose a heterogeneous-width pill blueprint for varied FL poisoning attacks.
+Unlike SRA’s one-time injection, our approach gradually poisons the global model during training, improving robustness against various defenses.
+2.3 Defenses against Poisoning Attacks in FL
+Existing defenses can be categorized based on the mitigation strategies that they utilize, including Adaptive Client Filtering, Statistical Parameter Aggregation, Client-dominant Detection, and Other Advanced Metrics and Pipelines.
+To comprhensively evaluate our method, we use Multi-Krum (MKrum) (krum), Trimmed Mean (Trim) (yin2018trim), Coordinate-wise Median (Median) (yin2018trim), Bulyan (bulyan), FLTrust (cao2020fltrust), FLDetector (FLD) (zhang2022fldetector), DnC (shejwalkar2021manipulating), and Flame (nguyen2022flame), a set of representative defenses, as our baselines.
+More details are in the extended version.
+2.4 Threat Model
+We follow the typical threat model used in existing studies (fang2020local; shejwalkar2021manipulating), where the attacker has access to a subset of compromised clients and aims to increase the error rates of the global model on specific classes or across all classes.
+In this scenario, defenses cannot directly analyze the data on each client as the defender’s setting in  (krum; yin2018trim; cao2020fltrust; guo2021siren).
+Instead, they identify malicious clients by analyzing the uploaded client updates.
+Further details are in the extended version.
 Figure 1: Overview of our augmentation method. The red parts indicate our augmentation method’s contribution, and the cyan parts represent the standard federated learning architecture.
-2.3 Defenses against Poisoning Attacks in FL
-Existing defenses can be categorized based on the mitigation strategies that they utilize.
-Adaptive Client Filtering.
-These techniques such as Krum and Multi-Krum [20] filter out malicious clients through single or multiple rounds of client selection based on distance scores.
-FLTrust [21] computes trust scores using the cosine similarity between each client update and the server model’s update for weighted averaging.
-SignGuard [22] employs sign-based clustering combined with norm-based thresholding to identify and filter malicious clients.
-Flame [23] and Deepsight [24] propose adaptive clustering and clipping to safeguard against backdoor attacks.
-SkyMask [25] clusters trainable feature masks of clients to assess each client’s risk level.
-Statistical Parameter Aggregation.
-Approaches like Median and Trim [26] use coordinate-wise median or trimmed mean values to aggregate model updates.
-Bulyan [27] enhances robustness by integrating Krum with Trim techniques.
-Fool’s Gold [28] applies an adaptive learning rate based on inter-client contribution similarity to mitigate the effects of malicious updates.
-SparseFed [29] aggregates sparsified updates, reducing the risk of model poisoning attacks.
-Client-dominant Detection.
-Siren [30] and Siren+ [31] set proactive accuracy-based alarms at the client level with the corresponding server-side decisions to counter various model poisoning attacks.
-FL-WBC [32] introduces client-side noise to diminish the efficacy of attacks and shorten their duration.
-FLIP [33] achieves higher robustness through client-side reverse-engineering defenses against extensive poisoning strategies.
-LeadFL [34] uses a client-side Hessian matrix optimization to reduce the impact of adversarial patterns on backdoor and targeted attacks.
-Other Advanced Metrics and Pipelines.
-Various studies employ other sophisticated metric pipelines designed for detection to ensure robust defense against poisoning attacks.
-These include techniques proposed in studies such as Zeno [35], CRFL [36], FedRecover [37], FLCert [38], FLDetector [39], and MESAS [50].
-To comprhensively evaluate our method, we use Multi-Krum (MKrum) [20], Trimmed Mean (Trim) [26], Coordinate-wise Median (Median) [26], Bulyan [27], FLTrust [21], FLDetector (FLD) [39], DnC [6], and Flame [23], a set of representative methods, as our baselines.
-More details are shown in Appendix A.
-3 Threat Model
-Attacker’s Goal and Capabilities.
-This paper focuses on improving the effectiveness of existing poisoning attacks in FL.
-Similar to previous work [4, 6], an attacker aims to raise the error rates of the global model on a specific class or multiple classes by sending poisoned model updates via compromised clients during the iterative aggregation.
-Our method does not require any additional knowledge compared with existing FL poisoning attack.
-Hence we reuse the typical threat model in existing studies [4, 6].
-The attacker has a complete control of the compromised clients, including their local data, local training, and uploading process.
-With the aggregated resources on the compromised clients, the attacker may aggregate the local data from the compromised clients to do extra training or aggregate their local updates to do model estimation.
-The attacker may or may not know the updates of other benign clients, depending on the confidentiality of the communication channels between the server and clients.
-Besides, the attacker cannot access the server’s information, including the aggregation rules or the selected clients in each round.
-Defense Settings.
-Most of the defenses in FL are deployed and executed on the server.
-We adopt a similar defense setting as existing studies [20, 26, 21, 30].
-The server cannot directly analyze the local data or the local training of clients.
-It can only detect malicious clients through model updates from different clients.
-The server can collect and possess a root test dataset to provide more accurate and robust detection, while the data of such a root test dataset cannot be derived from clients.
-The data distribution of this root test dataset may or may not be the same as the data distribution across the clients.
 3 Design Objectives and Challenges
@@
 Corresponding to each objective, three challenges need to be addressed:
-  • It presents a significant challenge that the attack augmentation method must use significantly fewer parameters while still achieving similar results as the original attacks.
+  • It presents a significant challenge that the attack augmentation method must use significantly fewer parameters while still achieving similar attack effects.
   • It is challenging to develop a uniform augmentation method for various FL poisoning attacks since they require different information and are implemented in different training stages.
@@
 4 Design
-TABLE I: Main notations. Symbols in the gray part are used for attacks.
+4.1 Overview of Our Method
+We are the first that propose a universal attack augmentation pipeline for most FL poisoning attacks, considering all of the stealthiness, compatibility, and generality.
+Figure 1 presents the three key stages.
+In the extended version, we include a comprehensive table of all main symbol notations used in the paper, along with an algorithm table detailing the full workflow of our method.
+[…]
+[…]
+[…]
+4.2 Pill Construction
+This stage aims to construct a pill structure for augmenting the stealthiness while retaining the original attack effectiveness.
+The pill is carefully crafted to only involve a minimal subset of parameters from specific positions of the target model.
+We first define a pill’s blueprint as the pill’s graphic structure, independent of target model parameters.
+Then, we propose a dynamic pill search algorithm to identify and map concrete parameters from the target model to the blueprint.
+Designing Pill Blueprint.
+Inspired by SRA (qi2022sra), which demonstrates that poisoning a narrow subnetwork (one neuron/channel per layer) can effectively implant backdoors in machine learning models (outside the FL setting), we adapt and generalize the concept for FL.
+The original SRA design is unsuitable for our goals as it employs a fixed, pre-selected subnet tailored to a specific architecture, ignores FL’s training dynamics, lacks adaptability to different targets, and produces non-stealthy subnets with disproportionately large weights to propagate the poison through a small network.
+We introduce a novel blueprint method with a general subnet structure whose instantiations vary dynamically during FL training.
+A search algorithm selects important neurons at each step, enabling small weight modifications while ensuring effective dissemination and high stealthiness.
+The blueprint also supports multiple targets by simultaneously manipulating outputs across all relevant neurons.
+Specifically, our pill blueprint design follows the rules below (more details shown in the extended version):
+  1. 1.
+    The pill blueprint only contains one neuron in each linear layer or one channel in each convolutional layer, except for the last two layers.
+Suppose $\mathcal{N}_{i}^{p}$ represents the neuron/channel number in Layer $i$ in our pill blueprint, then $\mathcal{N}_{i}^{p}=1$ when i<L1i<L-1, where $L$ is the total layer counts in our pill blueprint.
+  2. 2.
+    In the last two layers of our pill blueprint, $\mathcal{N}^{p}_{L-1}=\mathcal{N}^{p}_{L}=$ number of classes.
+Figure 2: Cosine similarity between FLTrust server and malicious updates with different extra local epochs.
+[…]
+Figure 3: Intuition behind distance-based adjustment in our augmentation method.
+Dynamic Pill Search.
+According to existing studies on neural network pruning (frankle2018lottery; lin2018deep; han2015deep; mugunthan2022fedltn; jiang2022model), parameters with a larger magnitude typically dominate the model’s performance.
+Thus, the best approach is to analyze the parameters to find a globally optimal pill that includes the most critical parameters.
+However, such a globally optimal pill could be identified via a pruning-based method (wu2020mitigating; sun2021soteria), and hence our attack could be easily detected.
+Besides, searching for a globally optimal pill is inefficient when the model has a large number of parameters.
+Thus, we search for an approximate pill instead, with an attacker-defined start point, and only evaluate a small subset of the entire model’s parameters.
+We name the search algorithm as “approximate max pill search”.
+The key idea is to perform a targeted neuron search at each layer by focusing only on the neurons connected to the selected neurons from the previous layer, following a high-sum-of-weights-first principle that prioritizes neurons based on the cumulative sum of their connection weights to the previously selected neurons.
+The entire search contains four steps:
+Step 1.
+Random Start Point Selection: Randomly select neurons from the first layer of the target model as start points, denoted as $\mathcal{V}_{1}$, based on the neuron count $\mathcal{N}^{p}_{1}$ in pill blueprint’s first layer.
+Such selected start point neurons are fixed throughout the search.
+Step 2.
+Layer-wise Search: For each subsequent layer $l_{i}$, we compute the sum of weights connecting neurons in $\mathcal{V}_{i-1}$ to neurons in $l_{i}$ and rank the neurons in $l_{i}$ based on the descending order of the sum of weights.
+Top $\mathcal{N}^{p}_{i}$ neurons are chosen for $\mathcal{V}_{i}$, where $\mathcal{N}^{p}_{i}$ denotes the number of neurons in the pill blueprint’s $i$-th layer.
+Step 3. Output Neuron Pairing: Pair the selected neurons $\mathcal{V}_{L-1}$ in the final hidden layer with the neurons in the output layer $l_{L}$, ensuring a one-to-one correspondence.
+Step 4.
+Pill Mask Construction: Two masks are constructed.
+$M$ marks the pill parameters in the target model, and $M_{disc}$ records the disconnection locations between the pill and the remained neurons in the target model.
+The searched pill ensures both effectiveness and stealthiness in attacks.
+Detailed information for each step is provided in the extended version, along with a concrete example and an overhead analysis of the pill search algorithm.
+4.3 Pill Poisoning
+In the pill poisoning stage, we aim to condense the poison into the pill using existing attacks.
+To achieve compatibility, our method simply reuses existing FL poisoning attacks, without any intrusive modification to their original implementations.
+We only modify the input of existing FL poisoning attacks by replacing the base model update with the update from a model that has undergone extra training rounds, denoted as $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$.
+Additionally, we restrict changes to parameters within the pill.
+The output is a poisoned pill that will be used in the next pill injection stage.
+The motivation to use an extra-trained model update as the reference model update is shown in Figure 3.
+As shown in the figure, with the increasing number of extra training rounds on the malicious clients, the generated malicious model update becomes less opposite to the FLTrust (cao2020fltrust) server’s model update.
+Thus, we adopt the extra training in our method and limit the extra training epoch number $E_{extra}$ to less than the number of malicious clients $m$ times the benign local training epoch number $E$, denoted as $E_{extra}\leq m\cdot E$.
+This constraint ensures compliance with the threat model, as the attacker can utilize the data and computational resources of compromised clients.
+4.4 Pill Injection
+In the pill injection stage, we aim to inject the pill into the model and use a two-step adjustment method to further camouflage the pill.
+Thus, the entire injection stage could be divided into two parts - pill injection and camouflaging.
+After this stage, the poison pill is seamlessly integrated with the benign model update and uploaded to the FL server.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
+Sign-Flipping 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
+[…]
+Trim Attack 0.243 0.109 0.139 0.146 0.174 0.179 0.116 0.332 0.120 0.201 0.163 0.231 0.238 0.124
+[…]
+Krum Attack 0.116 0.109 0.189 0.201 0.172 0.137 0.786 0.128 0.116 0.235 0.276 0.217 0.160 0.947
+[…]
+Min-Max Attack 0.183 0.110 0.431 0.330 0.183 0.218 0.825 0.269 0.125 0.619 0.434 0.255 0.278 0.831
+[…]
+Table 1: Error rates on Fashion-MNIST under cross-silo setting with 20% malicious clients in the 50-client FL system.
+Pill Insertion & Disconnection.
+In this part, our goal is to insert the pill into the model, and minimize the impact of the benign model updates on our pill.
+We use an estimated global model update as the benign model update, which is estimated as the coordinate-wise mean values of all the normal model updates from the compromised clients.
+The estimation process is hence presented as Equations (2) (Estimation() in Algorithm in the extended version),
+$\Delta\boldsymbol{\widetilde{g}}_{t+1}$ $\leftarrow\mathop{mean}\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\}$, (2)
+where $\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}$ is the normal updates from the compromised clients.
+By aggregating information from multiple malicious clients, the estimated global model update is more similar to the genuine one, providing more budget for our poison pill.
+After obtaining the estimated global model update $\boldsymbol{\Delta\widetilde{g}}_{t+1}$, we directly replace the parameters corresponding to the pill parameters (which have been poisoned in the previous stage) via the pill’s mask $\boldsymbol{M}$.
+Then, we replace the parameters that connect the pill and the other estimated global model updates with the disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$, using the disconnection mask $\boldsymbol{M}_{disc}$.
+The disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$ is calculated as $0-\boldsymbol{g}_{t}$, and is bounded by the maximum and minimum values of the reference model update $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$.
+The disconnection update gradually reduces the connection parameters between the pill and the rest of the model to 0, and finally isolates the poison pill from the global model, guaranteeing the attacking effects.
+Pill Adjustment.
+After the injection, we use a two-step adjustment to further adjust the pill, improving the generality against multiple detection metrics simultaneously.
+In this stage, we consider two prevailing detection metrics – distance and cosine similarity.
+To increase the cosine similarity between the poisoned model update and the benign model update in our method, we balance the magnitudes of both the poison pill’s parameters and the other benign parameters.
+Similarly, to minimize the distance discrepancy between the poisoned and benign model updates, we adjust the magnitude of the entire poisoned model update, as shown in Figure 3.
+Thus, we first use the similarity-based adjustment, then use the distance-based adjustment, balancing the effectiveness and the stealthiness of the poisoned model update.
+This two-step adjustment is particularly effective when combined with our method, which selectively poisons only a tiny subset of the model’s parameters.
+By altering just a few parameters, our method preserves a substantial number of benign parameters, which are crucial for making effective adjustments.
+As a result, the poisoned model update can bypass a wide range of defenses since they are typically designed based on the combination or variants of distance and cosine similarity metrics, and they usually do not anticipate such a focused and minimal interference in the model parameters.
+More details are shown in the extended version.
+5 Evaluation
+This section evaluates how our method enhances existing FL poisoning attacks from three perspectives.
+First, we assess its Augmentation Effectiveness against four poisoning attacks using nine state-of-the-art defenses across three datasets (Section 5.2).
+Second, we examine its Stealthiness under two detection metrics (Section 5.3).
+Finally, we conduct a Generality Analysis, testing different malicious client proportions, both cross-silo and cross-device settings, and various pill search rules (Section 5.4).
+Our method substantially strengthens existing attacks, bypassing all nine defenses in over 90% of cases and raising error rates by up to seven times compared to the originals.
+It remains effective across diverse data distributions, model architectures, malicious client proportions, and pill search rules.
+5.1 Evaluation Settings
+In our experiments, the malicious client proportion is set to $20\%$ by default.
+We assess 9 standard aggregation rules, including FedAvg, FLTrust, Multi-Krum, Bulyan, Median, Trim, FLDetector, DnC, and Flame, alongside 4 notable model poisoning attacks: sign-flipping, Trim attack, Krum attack, and Min-Max.
+Experiments utilize 50 clients for MNIST (lecun1998mnist) and Fashion-MNIST (fashionmnist), and 30 for CIFAR-10 (cifar), accommodating cross-silo and cross-device scenarios.
+Implementation is done in PyTorch (paszke2019pytorch).
+More details are provided in the extended version.
+Figure 4: Comparison of Multi-Krum distance score between benign updates and malicious updates when using original poisoning attacks with and without our method.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.106 0.104 0.103 0.108 0.127 0.107 0.116 0.111 0.119 0.113 0.113 0.140 0.114 0.123
+Sign-Flipping 0.964 0.109 0.108 0.110 0.130 0.108 0.117 0.909 0.119 0.114 0.119 0.144 0.120 0.125
+[…]
+Trim Attack 0.112 0.111 0.111 0.115 0.132 0.114 0.116 0.125 0.115 0.121 0.125 0.153 0.122 0.122
+[…]
+Krum Attack 0.107 0.108 0.114 0.123 0.141 0.112 0.668 0.116 0.117 0.124 0.138 0.173 0.122 0.410
+[…]
+Min-Max Attack 0.117 0.108 0.118 0.135 0.142 0.128 0.111 0.124 0.119 0.142 0.166 0.162 0.145 0.136
+[…]
+Table 2: Error rates on Fashion-MNIST under cross-silo setting with 10% malicious clients in the 50-client FL system.
+Distribution IID Non-IID
+Attack FAvg FLT MKr Bulyan Med Trim DnC FLD Flame FAvg FLT MKr Bulyan Med Trim DnC FLD Flame
+No Attack 0.48 0.48 0.50 0.46 0.55 0.45 0.44 0.49 0.49 0.48 0.47 0.49 0.49 0.58 0.52 0.46 0.50 0.53
+Sign-Flipping 0.89 0.47 0.58 0.53 0.62 0.46 0.46 0.49 0.50 0.90 0.51 0.51 0.62 0.65 0.57 0.50 0.60 0.53
+[…]
+Trim ATK 0.48 0.50 0.48 0.53 0.62 0.51 0.45 0.45 0.50 0.57 0.49 0.60 0.59 0.65 0.54 0.48 0.48 0.50
+[…]
+Krum ATK 0.47 0.54 0.47 0.56 0.54 0.51 0.45 0.80 0.50 0.48 0.50 0.49 0.52 0.64 0.51 0.48 0.89 0.50
+[…]
+Min-Max ATK 0.45 0.50 0.46 0.50 0.57 0.46 0.51 0.52 0.52 0.47 0.50 0.49 0.56 0.63 0.60 0.47 0.48 0.48
+[…]
+Table 3: Error rates on CIFAR-10 under cross-silo setting with 20% malicious clients in the 30-client FL system.
+5.2 Augmentation Effectiveness
+This section provides a detailed analysis of our method’s augmentation effect on Fashion-MNIST in a 50-client cross-silo FL system with 20% malicious clients.
+We test our method on both IID and non-IID data, resulting in an average error rate increase of over 0.25 for all baseline attacks, showing our method’s effectiveness and compatibility.
+Results on IID Data.
+The error rates of four baseline FL poisoning attacks, with and without our method, are shown in the left half of Table 1.
+Our method enhances the error rates of the existing poisoning attacks in $23$ out of $28$ scenarios, against FedAvg and five defenses.
+The maximum increase in error rate is $0.658$, and the average increase reaches $0.274$.
+This substantial elevation from the attack-free baseline error rate of $0.109$ underscores our method’s capability to significantly compromise existing defenses’ integrity.
+Results on Non-IID Data.
+Evaluations on non-IID data further validate the effectiveness of our method, demonstrating its superiority in $23$ of $28$ cases.
+The highest error rate increase reaches $0.637$, with an average increase of $0.281$.
+Although there is a slight reduction in the maximal error rate increase in the non-IID setting, these results still demonstrate our method’s ability to effectively enhance attacks in more complex and heterogeneous data environments.
+All attacks augmented by our method can bypass all baseline defenses, including FLTrust and FLDetector, with the exception of the sign-flipping attack.
+Notably, the Min-Max attack demonstrates superior effectiveness in non-IID data settings, achieving significant improvements compared to its performance on IID data.
+Other attacks also exhibit similar error rate improvements relative to their results on IID data, indicating that our method maintains its robustness and effectiveness in more complex data environments.
+More detailed analyses are presented in the extended version.
+5.3 Stealthiness Analysis
+To further assess our method, we examine its stealthiness during FL training, focusing on its impact on the distance and cosine similarity scores of existing poisoning attacks.
+Results show that our method can make malicious clients appear as benign, or even more “benign” than genuine clients, owing to its pill design with distance-based and similarity-based adjustments.
+As shown in Figure 4, the average distance scores of malicious clients (with our method) across four baseline model poisoning attacks closely match or even coincide with those of benign clients throughout training.
+Detailed results, including similarity score analyses, are provided in the extended version.
+5.4 Generality Analysis
+In this section, we further discuss the generality of our method across three key factors: malicious client proportion, client participation frequency, and datasets & model architectures.
+The results indicate that our method consistently maintains its effectiveness despite changes in these conditions, demonstrating its adaptability and broad applicability in augmenting existing attacks.
+Impact of The Malicious Client’s Proportion.
+We first assess the effectiveness of our method in both IID and non-IID cross-silo FL systems with only 1010% of clients compromised, as shown in Table 2.
+This setup reveals that all baseline model poisoning attacks yield lower error rates on the global model compared with scenarios with 2020% compromised clients.
+While the increase in error rates is less than those in the 2020% compromised client scenario, our method still effectively raises the global model’s error rates in 2525/$26$ out of $28$ cases (IID/non-IID setting).
+The maximum increase in error rates reaches $0.403$, with an average increase of $0.144$.
+This average is notably higher (¿22x higher) than the error rates observed in attack-free FL conditions.
+Specifically, our method helps sign-flipping/Trim/Krum/Min-Max attacks achieve an average error rate increase of 0.1330.133/0.0940.094/0.1360.136/$0.209$.
+More detailed results are presented in the extended version.
+Impact of The Client Participation Frequency.
+We then extend the evaluation of our method to a cross-device FL system, where only 4040% of clients are selected for participation in each communication round.
+This setup results in less frequent participation from each client and a fluctuating proportion of malicious clients across different rounds.
+The maximum error rate increase with our method is $0.639$, with an average increase across different attacks and defenses of $0.279$.
+These results are consistent with those from the cross-silo FL system, underscoring our method’s effectiveness and generality across different FL configurations.
+This evaluation demonstrates our method’s robust performance and adaptability, not only in a controlled cross-silo environment but also under the more various conditions in cross-device FL systems.
+More details are presented in the extended version.
+Impact of The Datasets and Model Architectures.
+Following the evaluation with the Fashion-MNIST dataset, we test our method on the MNIST and CIFAR-10 dataset, employing the four-layer CNN model and the AlexNet model to further verify our method’s generality across different datasets.
+The collective results show that our method performs even better with larger datasets or more complex machine learning models.
+This trend confirms the generality of our method by revealing its capability to maintain consistent performance enhancements regardless of the dataset or model complexity involved.
+Specifically, our method helps all four baseline attacks bypass all nine baselines on CIFAR-10 dataset, achieving $0.288$ error rate increase on average, presented in Table 3.
+More detailed results on MNIST dataset are shown in the extended version.
+Beyond the three key factors, we also investigate the impact of the pill search algorithm in the extended version.
+The results indicate that the “approximate max pill search” algorithm outperforms the “approximate min pill search” in 41 out of 56 cases ($\sim$73%), highlighting its effectiveness in selecting the most influential parameters to maximize attack impact.
+Additional findings on ablation studies and generalizability are also provided in the extended version.
+6 Discussion
+To further evaluate the robustness of our method when defenses are aware of the attack strategies (white-box scenario), we design an adaptive defense and present the experimental details in the extended version.
+Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics, it remains insufficient to thwart the enhanced capabilities of our method.
+We also presented a detailed discussion of the limitations and future directions in the extended version.
+7 Conclusion
+We propose a novel attack-agnostic augmentation method to enhance FL poisoning attacks by concentrating them into a tiny subnet (i.e., pill).
+Our approach comprises pill construction, pill poisoning, and pill injection, enabling existing FL poisoning attacks to achieve over $2\times$ higher error rates on average.
+By intensifying the inherent vulnerabilities in current FL defenses, our method underscores the urgent need for more robust and fine-grained detection mechanisms.
+Acknowledgements
+We thank the anonymous reviewers for their valuable feedback.
+This work was supported in part by the National Natural Science Foundation of China (NO.
+62472284, 62572426), Shanghai Key Laboratory of Scalable Computing and Systems.
+The work of H.
+Wang was supported in part by the United States National Science Foundation (NSF) under grants 2534286, 2523997, 2315612, and 2332638.
+Appendix A Additional Details of Existing Attacks and Defenses
+A.1 Additional Details of Existing Attacks in FL
+Although model poisoning attacks are effective, existing attacks have limited stealthiness and can be detected by many existing defenses.
+Our goal is hence to demonstrate that such attacks can be augmented in a uniform way.
+Model poisoning attacks directly manipulate the parameters uploaded by clients, with a minimal interference to the local training process.
+Among these attacks, the simplest form is the sign-flipping attack, which directly flips the model update and scales it with a constant factor.
+A-Little-is-Enough (baruch2019little) generates malicious updates within a calculated perturbation range to deceive the global model.
+Adaptive attacks (fang2020local), such as the Trim attack and the Krum attack, dynamically scale malicious updates based on parameter values and distances.
+The Min-Max and Min-Sum Attacks (shejwalkar2021manipulating) provide a dynamic scaling for malicious updates based on different distance-based criteria.
+MPAF (cao2022mpaf) aims to drive the global model towards a predefined target model with poor performance on given FL tasks.
+For the sake of generality, we employ the sign-flipping attack, two types of adaptive attacks (fang2020local) (the Trim attack and the Krum attack), and the Min-Max attack (shejwalkar2021manipulating) as the baseline attacks in this paper.
+A.2 Additional Details of Existing Defenses in FL
+Adaptive Client Filtering.
+These techniques such as Krum and Multi-Krum (krum) filter out malicious clients through single or multiple rounds of client selection based on distance scores.
+FLTrust (cao2020fltrust) computes trust scores using the cosine similarity between each client update and the server model’s update for weighted averaging.
+SignGuard (xu2021signguard) employs sign-based clustering combined with norm-based thresholding to identify and filter malicious clients.
+Flame (nguyen2022flame) and Deepsight (rieger2022deepsight) propose adaptive clustering and clipping to safeguard against backdoor attacks.
+SkyMask (yan2023skymask) clusters trainable feature masks of clients to assess each client’s risk level.
+Statistical Parameter Aggregation.
+Approaches like Median and Trim (yin2018trim) use coordinate-wise median or trimmed mean values to aggregate model updates.
+Bulyan (bulyan) enhances robustness by integrating Krum with Trim techniques.
+Fool’s Gold (foolsgold) applies an adaptive learning rate based on inter-client contribution similarity to mitigate the effects of malicious updates.
+SparseFed (panda2022sparsefed) aggregates sparsified updates, reducing the risk of model poisoning attacks.
+Client-dominant Detection.
+Siren (guo2021siren) and Siren+ (guo2024siren+) set proactive accuracy-based alarms at the client level with the corresponding server-side decisions to counter various model poisoning attacks.
+FL-WBC (sun2021flwbc) introduces client-side noise to diminish the efficacy of attacks and shorten their duration.
+FLIP (zhang2023flip) achieves higher robustness through client-side reverse-engineering defenses against extensive poisoning strategies.
+LeadFL (zhu2023leadfl) uses a client-side Hessian matrix optimization to reduce the impact of adversarial patterns on backdoor and targeted attacks.
+Other Advanced Metrics and Pipelines.
+Various studies employ other sophisticated metric pipelines designed for detection to ensure robust defense against poisoning attacks.
+These include techniques proposed in studies such as Zeno (xie2019zeno), CRFL (xie2021crfl), FedRecover (cao2023fedrecover), FLCert (cao2022flcert), FLDetector (zhang2022fldetector), and MESAS (krauss2023mesas).
+Here are more details of the baseline defenses used in our paper:
+Krum and Multi-Krum (MKrum) (krum).
+Krum uses a distance score as the metric.
+In each round, the Krum server sums the distances between each client update $\boldsymbol{g}_{t}^{(i)}$ and its Km2K-m-2 neighbors, and uses these sums as the scores for all the clients.
+The Krum server then selects the client’s model update with the lowest score.
+Multi-Krum is a variant of Krum that uses iterative Krum to pick multiple candidates for aggregation.
+Coordinate-wise Median (Median) (yin2018trim).
+Coordinate-wise Median (Median) uses the per-parameter median values of the model updates from the clients as the aggregated global model update, which is then used to generate the next-round global model.
+Trimmed Mean (Trim) (yin2018trim). Trimmed Mean (Trim) calculates per-parameter trimmed mean values of the client model updates and packs them as the global model update.
+Bulyan (bulyan).
+Bulyan is a combination of Krum and Trim.
+It first uses the Krum-based method to select multiple candidates, and uses the per-parameter trimmed mean values of the candidate model updates as the final global model update.
+FLTrust (cao2020fltrust).
+FLTrust trains a server model with a small root dataset.
+In each round, it computes the clipped cosine similarities between the server model update and client updates as trust scores, and then uses the trust scores as weights to aggregate all the normalized client model updates.
+FLDetector (FLD) (zhang2022fldetector).
+FLDetector filters out malicious clients by checking the multi-round consistency of all client updates.
+Malicious updates typically have lower consistency compared to benign ones.
+Flame (nguyen2022flame).
+Flame utilizes HDBSCAN-based (hdbscan) dynamic clustering to filter out malicious clients, and aggregates median-clipped benign updates with adaptive noise as the global model update.
+Layer Type       CNN ModelOriginal       Pill BlueprintOur
+Input $\times 28\times 1 28×28$$\times 28\times 1$
+Conv2d $\times 3\times 30 3×3$$\times 3\times 1$
+ReLU -1 -
+MaxPool2d $\times 2 2×2$$\times 2$
+Conv2d $\times 3\times 50 3×3$$\times 3\times 1$
+ReLU - -
+MaxPool2d $\times 2 2×2$$\times 2$
+Linear $\times 100 25×1025\times 10$
+ReLU - -
+Linaer $0\times 10 1010$
+Softmax - $\times{\textsuperscript{2}}$
+  • 1
+    “-” represents that the model has this layer with no specified configuration.
+  • 2
+    $\times$” represents that the model does not contain this layer.
+Table 4: Architectures of the original CNN model and the corresponding pill blueprint.
+Layer Type      AlexNetOriginal       Pill BlueprintOur
+Input $\times 32\times 3 32×32$$\times 32\times 3$
+Conv2d $\times 11\times 48 11×11$$\times 11\times 1$
+ReLU - -
+MaxPool2d $\times 3 3×3$$\times 3$
+Conv2d $\times 3\times 96 3×3$$\times 3\times 1$
+ReLU - -
+MaxPool2d $\times 3 3×3$$\times 3$
+Conv2d $\times 3\times 192 3×3$$\times 3\times 1$
+ReLU - -
+Conv2d $\times 3\times 192 3×3$$\times 3\times 1$
+ReLU - -
+Conv2d $\times 3\times 128 3×3$$\times 3\times 1$
+ReLU - -
+MaxPool2d $\times 3 3×3$$\times 3$
+Linear $4608\times 1024$ $36\times 1$
+ReLU - -
+Linear $024\times 512 1×1$$\times 10$
+ReLU - -
+Linear $512\times 10$ $10$
+Softmax - $\times$
+  • 1
+    “-” represents that the model has this layer with no specified configuration.
+  • 2
+    $\times$” represents that the model does not contain this layer.
+Table 5: Example architectures of original AlexNet and the corresponding pill blueprint.
 Symbol Meaning
@@
 $\boldsymbol{g}_{t}$ Global model in round $t$
-$lr$ Learning rate
+lrlr Learning rate
 $f()$ Loss function used in the FL training
@@
 $C_{\downarrow}$ Down-scaling factor
-1 function MalUpdate($i$,$t$,$\boldsymbol{g}_{t}$)
+Table 6: Main notations. Symbols in the gray part are used for attacks.
+Appendix B Detailed Threat Model
+Attacker’s Goal and Capabilities.
+This paper focuses on improving the effectiveness of existing poisoning attacks in FL.
+Similar to previous work (fang2020local; shejwalkar2021manipulating), an attacker aims to raise the error rates of the global model on a specific class or multiple classes by sending poisoned model updates via compromised clients during the iterative aggregation.
+Our method does not require any additional knowledge compared with existing FL poisoning attacks.
+Hence we reuse the typical threat model in existing studies (fang2020local; shejwalkar2021manipulating).
+The attacker has a complete control of the compromised clients, including their local data, local training, and uploading process.
+With the aggregated resources on the compromised clients, the attacker may aggregate the local data from the compromised clients to do extra training or aggregate their local updates to do model estimation.
+The attacker may or may not know the updates of other benign clients, depending on the confidentiality of the communication channels between the server and clients.
+Besides, the attacker cannot access the server’s information, including the aggregation rules or the selected clients in each round.
+Defense Settings.
+Most of the defenses in FL are deployed and executed on the server.
+We adopt a similar defense setting as existing studies (krum; yin2018trim; cao2020fltrust; guo2021siren).
+The server cannot directly analyze the local data or the local training of clients.
+It can only detect malicious clients through model updates from different clients.
+The server can collect and possess a root test dataset to provide more accurate and robust detection, while the data of such a root test dataset cannot be derived from clients.
+The data distribution of this root test dataset may or may not be the same as the data distribution across the clients.
+[…]
  % 1.Pill Construction
@@
 8 $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}\leftarrow\boldsymbol{g}_{t}-\boldsymbol{\hat{g}}^{m}_{t+1}$;
-9       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$ Poisoning($i$,$t$,$\boldsymbol{param}$,$\Delta\boldsymbol{\hat{g}}^{m}_{t+1}$);
+[…]
 10 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
-       % 3. Poison Pill Injection
+ % 3.Pill Injection
 11 $\mathop{param}\leftarrow\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\}$;
-12       $\Delta\boldsymbol{\widetilde{g}}_{t+1}\leftarrow$ Estimation($i$,$t$,$\boldsymbol{g}_{t}$,$\mathop{param}$);
+[…]
 13 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow\Delta\boldsymbol{g}_{t+1}^{(i)}+(\mathbf{1}-\boldsymbol{M})\odot\boldsymbol{\Delta\widetilde{g}}_{t+1}$;
@@
 16 $\mathop{param}=\mathop{param}\bigcup\{\boldsymbol{M}_{all}=M+M_{disc}\}$;
-17       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
-18       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
+17 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrowSimAdjust$($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
+18 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrowDistAdjust$($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
 19 return $\Delta\boldsymbol{g}^{(i)}_{t+1}$
-Algorithm 1 Our augmentation method’s workflow. Without losing generality, we assume the first $m$ clients as malicious clients.
-5.1 Overview of Our Method
-Fig. 1 presents the three key stages. Table I presents notations of main symbols in this paper.
-[…]
-[…]
-[…]
-We are the first that propose a universal attack augmentation pipeline for most FL poisoning attacks, considering stealthiness, compatibility, and generality.
-The detailed workflow of our method is shown in Algorithm 1.
-5.2 Pill Construction
+Algorithm 1 Our method’s workflow. Assume the first $m$ clients as malicious clients.
+Appendix C Main Notations and Pseudo-Code
+Table 6 presents the main notation used in the paper,W and Algorithm 1 shows the pseudo-code of our proposed attack augmentation method.
+Appendix D Additional Details of Concrete Pill Blueprints
+Table 4 and Table 5 illustrate the model structures of the CNN model and the simplified AlexNet, with their corresponding pill’s blueprints.
+Appendix E Detailed Pill Search Algorithm
 Figure 5: An example of the “approximate max pill search” algorithm in our augmentation method.
-This stage aims to construct a pill structure for augmenting the stealthiness while retaining the attack effectiveness before being augmented.
-The pill is carefully crafted to involve a minimal subset of parameters from specific positions of the target model.
-We first define a pill’s blueprint as the pill’s graphic structure, independent of target model parameters.
-Then, we propose a dynamic pill search algorithm to identify and map concrete parameters from the target model to the blueprint.
-Designing Pill Blueprint.
-The blueprint design is inspired by SRA [15], which shows that poisoning a narrow subnetwork (one neuron/channel in each layer) is adequate to effectively inject backdoors into machine learning models (not in the FL setting).
-However, their technique cannot be used for our purposes as their subnet architecture is very specific.
-It does not support attacking various targets; it is a fixed and pre-selected subnet without considering the dynamics of model training in FL; and its poisoned subnet is not stealthy, having substantially larger weight values compared to others due to the need to disseminate the poison effect through such a small pre-selected network.
-Therefore, we propose a novel blueprint method, in which the subnet structure is general, and its instantiations (i.e., the concrete subnets) vary across steps in the FL training procedure, including important neurons by a dynamic search algorithm.
-This allows small weight changes because poisoning important neurons enables easy dissemination, maximizing attack stealthiness.
-In particular, the pill blueprint is designed to accommodate various target classes of different FL poisoning attacks.
-It achieves this by manipulating the outputs relevant to multiple classes simultaneously, via disrupting all the output neurons together.
-Hence, our pill blueprint design follows the rules below:
-  1. 1.
-    The pill blueprint only contains one neuron in each linear layer or one channel in each convolutional layer, except for the last two layers.
-Suppose $\mathcal{N}_{i}^{p}$ represents the neuron/channel number in Layer $i$ in our pill blueprint, then $\mathcal{N}_{i}^{p}=1$ when $i<L-1$, where $L$ is the total layer counts in our pill blueprint.
-  2. 2.
-    In the last two layers of our pill blueprint, $\mathcal{N}^{p}_{L-1}=\mathcal{N}^{p}_{L}=$ number of classes.
-Dynamic Pill Search.
-According to existing studies on neural network pruning [40, 41, 42, 43, 44], parameters with a larger magnitude typically dominate the model’s performance.
-The optimal solution is hence to examine the model parameters to search for a globally optimal pill that encompasses the most important parameters.
-However, such a globally optimal pill could be identified via a pruning-based method [51, 52], and hence our attack could be easily detected.
-Besides, searching for a globally optimal pill is inefficient when the model has a large number of parameters.
-Thus, we search for an approximate pill instead, with an attacker-defined start point, and only evaluate a small subset of the entire model’s parameters.
-We name the search algorithm as “approximate max pill search”.
-A complete procedure of this search algorithm consists of the following four steps.
+A complete procedure of the “approximate max pill search” algorithm consists of the following four steps.
 To improve readability, we use “neuron” to represent both neurons in fully connected layers and channels in convolutional layers, and we use the classification task as an example:
@@
 Step 3 Output Neuron Pairing:
-After visiting $l_{L-1}$ layers, where $L$ is the total number of layers in the target model, $||\mathcal{V}_{L-1}||$ should equal to the neuron number in $l_{L}$ of the target model, which also equals to the number of classes.
+After visiting $l_{L-1}$ layers, where $L$ is the total number of layers in the target model, $\mathcal{V}_{L-1}||$ should equal to the neuron number in $l_{L}$ of the target model, which also equals to the number of classes.
 We select all the neurons in the target model’s $l_{L}$ layer into our pill to construct $\mathcal{V}_{L}$.
 Then, we only record the parameters from one neuron in $\mathcal{V}_{L-1}$ to only one neuron in $\mathcal{V}_{L}$ based on the index order (i.e., the first neuron in $\mathcal{V}_{L-1}$ is paired with the first neuron in $\mathcal{V}_{L}$).
-Since $||\mathcal{V}_{L-1}||$ equals to $||\mathcal{V}_{L}||$, the number of recorded parameters equals to the number of classes, avoiding poisoning too many parameters in a single layer.
+Since $\mathcal{V}_{L-1}||$ equals to $\mathcal{V}_{L}||$, the number of recorded parameters equals to the number of classes, avoiding poisoning too many parameters in a single layer.
 Step 4 Pill Mask Construction:
@@
 The two masks have the same shape as the target model’s parameters.
-To construct $\boldsymbol{M}$, we set the locations corresponding to the pill parameters to $1$, and the others to 00.
+To construct $\boldsymbol{M}$, we set the locations corresponding to the pill parameters to $1$, and the others to 0.
 Based on $\boldsymbol{M}$, we can similarly obtain the corresponding disconnected mask $\boldsymbol{M}_{disc}$, which sets the locations corresponding to the parameters from neurons other than $\mathcal{V}_{i-1}$ to $\mathcal{V}_{i}$ in each model layer $l_{i}$ (except for the $L$th layer since we choose all the neurons from it), and also those corresponding to parameters from $\mathcal{V}_{i-1}$ to other pill irrelevant neurons in each model layer $l_{i}$ to $1$.
@@
 Example.
-Fig.
-2 presents a concrete example of the search algorithm in a 4-layer linear model.
+Figure 5 presents a concrete example of the search algorithm in a 4-layer linear model.
 Since the start point is randomly selected by the attacker, defense methods can hardly guess it without any prior knowledge.
-In the example, suppose the model is a 4-layer linear models for a binary classification task.
+In the example, suppose the model is a 4-layer linear model for a binary classification task.
 Then the pill blue print contains one neuron in the first two layers, and contains two neurons in the last two layers.
 Initially, we randomly select a start neuron, specifically the second neuron in the first layer in the example.
-Then, we conduct layer-wise searching when visiting the second and third layer, selecting the parameters with the highest magnitudes.
+Then, we conduct layer-wise searching when visiting the second and third layers, selecting the parameters with the highest magnitudes.
 At the forth layer (output layer), we pair the two output neurons with the two selected neurons in the third layer based on the index order.
@@
 The computational complexity of our pill search is hence much smaller than the computational complexity of one round local training.
-TABLE II: Different dynamic patterns in our augmentation method, utilizing along with the max subnetwork searching.
 Notation Description
-$\textsc{Pattern}_{1}$ All layers use adaptive searching strategy
-$\textsc{Pattern}_{2}$ All layers use one-time searching strategy
-$\textsc{Pattern}_{3}$Centerstack[l]FE uses with adaptive searching strategy
-CLS uses repeated searching strategy
-$\textsc{Pattern}_{4}$Centerstack[l]FE uses repeated searching strategy
-CLS uses with adaptive searching strategy
-$\textsc{Pattern}_{5}$Centerstack[l]FE uses with adaptive searching strategy
-CLS uses with one-time searching strategy
-$\textsc{Pattern}_{6}$Centerstack[l]FE uses with one-time searching strategy
-CLS uses with adaptive searching strategy
-To make the pill search process dynamic, we also design several patterns to adaptively determine whether to change the pill in the training period, shown in Table II (FE represents the convolutional layers, CLS represents the linear layers).
-For more details about each specific dynamic pattern, please refer to Appendix C.
-The combination of the “approximate max pill search” with different dynamic patterns construct the complete dynamic pill search, considering both the stealthiness and the efficiency.
-Figure 3: Cosine similarities between FLTrust server’s model update and malicious model update when malicious clients use different extra training rounds.
-5.3 Pill Poisoning
-In the pill poisoning stage, we aim to condense the poison into the pill using existing attacks.
-To achieve compatibility, our method simply reuses existing FL poisoning attacks, without any intrusive modification to their original implementations.
-We only modify the input of existing FL poisoning attacks by replacing the base model update with the update from a model that has undergone extra training rounds, denoted as $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$.
-Additionally, we restrict changes to parameters within the pill.
-The output is a poisoned pill that will be used in the next pill injection stage.
-The motivation to use an extra-trained model update as the reference model update is shown in Fig.
-3.
-As shown in the figure, with the increasing number of extra training rounds on the malicious clients, the generated malicious model update becomes less opposite to the FLTrust [21] server’s model update.
-Thus, we adopt the extra training in our method and limit the extra training epoch number $E_{extra}$ to less than the number of malicious clients $m$ times the benign local training epoch number $E$, denoted as $E_{extra}\leq m\cdot E$.
-With this extra training epoch number limit, we do not violate the threat model since the attacker can utilize the data and computational resources of all the compromised clients.
-5.4 Pill Injection
-In the pill injection stage, we aim to inject the pill into the model and use a two-step adjustment method to further camouflage the pill.
-Thus, the entire injection stage could be divided into two parts - pill injection and camouflaging.
-After this stage, the poison pill is seamlessly integrated with the benign model update and uploaded to the FL server.
-Pill Insertion & Disconnection.
-In this part, our goal is to insert the pill into the model, and minimize the impact of the benign model updates on our pill.
-We use an estimated global model update as the benign model update, which is estimated as the coordinate-wise mean values of all the normal model updates from the compromised clients.
-The estimation process in the Estimation() in Algorithm 1 Line 11 is hence presented as Equations (2),
-$\Delta\boldsymbol{\widetilde{g}}_{t+1}$ $\leftarrow\mathop{mean}\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\},$ (2)
-where $\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}$ is the normal updates from the compromised clients.
-By aggregating information from multiple malicious clients, the estimated benign global model update is more similar to the genuine one, providing more budget for our poison pill.
-After obtaining the estimated global model update $\boldsymbol{\Delta\widetilde{g}}_{t+1}$, we directly replace the parameters corresponding to the pill parameters (which have been poisoned in the previous stage) via the pill’s mask $\boldsymbol{M}$.
-Then, we replace the parameters that connect the pill and the other estimated global model updates with the disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$, using the disconnection mask $\boldsymbol{M}_{disc}$.
-The disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$ is calculated as $0-\boldsymbol{g}_{t}$, and is bounded by the maximum and minimum values of the reference model update $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$.
-The disconnection update can gradually change the parameters of the connections between the pill and rest of the model to 00, and finally isolates the poison pill from the global model, guaranteeing the attacking effects of the poison pill.
-1 function SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
+$\textsc{Pattern}_{1}$ All layers use the adaptive searching strategy
+$\textsc{Pattern}_{2}$ All layers use the one-time searching strategy
+$\textsc{Pattern}_{3}$ CLS layers use the repeated searching strategyFE layers use the adaptive searching strategy
+$\textsc{Pattern}_{4}$ CLS layers use the adaptive searching strategyFE layers use repeated searching strategy
+$\textsc{Pattern}_{5}$ CLS layers use the one-time searching strategyFE layers use the adaptive searching strategy
+$\textsc{Pattern}_{6}$ CLS layers use the adaptive searching strategyFE layers use the one-time searching strategy
+Table 7: Different dynamic patterns in our method, utilizing along with “approximate max pill search”.
+To make the pill search process dynamic, we also design several patterns to adaptively determine whether to change the pill in the training period, shown in Table 7 (FE represents the convolutional layers, CLS represents the linear layers).
+For more details about each specific dynamic pattern, please refer to Appendix F.
+The combination of the “approximate max pill search” with different dynamic patterns constructs the complete dynamic pill search, considering both the stealthiness and the efficiency.
+Appendix F Additional Details of The Dynamic Patterns in Our Method
+We first design three searching strategies, including one-time searching strategy, repeated searching strategy, and adaptive searching strategy.
+In the one-time searching strategy, we search the pill based on the initial global model using the “approximate max pill search” algorithm introduced in § 4.2, and keep this pill unchanged in the whole FL training.
+The one-time searching strategy benefits the formation of pill in the global model, while the initial pill may be less effective with the increasing training rounds of the global model, due to the changing importance of the model parameters.
+On the contrary, the repeated searching strategy runs the ‘approximate max pill search” algorithm in every training round.
+The repeated searching strategy can help our method modify more parameters in the global model, and make the pill less traceable.
+While the attacking effects may be reduced due to the constantly changing pill.
+Considering advantages and disadvantages of both the one-time searching strategy and the repeated searching strategy, we design a more flexible searching strategy, termed as “adaptive searching strategy”.
+In the adaptive searching strategy, our method searches the new pill only when the pill is not successfully injected into the global model in the last round.
+The condition:
+$\texttt{Sim(}\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t},\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t}^{(i)}\texttt{)}<C_{search}$ (3)
+should be satisfied to trigger the new subnetwork searching on malicious client $i$, where $C_{search}$ is set as $0.94$ in the experiments.
+The adaptive searching strategy is a more moderate version of repeated searching.
+Since the three searching strategies have their unique advantages, we investigate different combinations of them in the experiments.
+We further divide the neural network into Feature Extractor (FE) and Classifier (CLS).
+Refer to the CNN model we used, the convolutional layers are regarded as FE, and the linear layers are regarded as CLS.
+We use different searching strategies in FE and CLS, respectively.
+In all the nine combinations, we test and keep six of them, noted as Pattern1 to Pattern6, shown in Table 7 in § 4.2.
+Such six patterns construct the entire dynamic pattern set used in our method.
+Appendix G Additional Details of Pill Adjustment
+1
+function SimAdjust($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
 2 $\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1},\boldsymbol{M}_{all}\}\leftarrow\mathop{param}$;
+3 $S_{max}\leftarrow\mathop{max}(0,\mathop{max}\{$Sim($\Delta\boldsymbol{\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)\prime}$$);i\in\{1,\cdots,m\}\}$);
+4 $\mathop{iter}\leftarrow 0$;
 […]
-4       $\mathop{iter}\leftarrow 0$;
-5       while Sim($\Delta\boldsymbol{\widetilde{g}}_{t+1},\Delta\boldsymbol{g}_{t+1}^{(i)}$)<$S_{max}$& $\mathop{iter}<C_{iter}$ do
 6    if $\mathop{iter}\%2$ then
@@
 16
-17
-18function DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
+17function DistAdjust($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
 18 $\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1},\boldsymbol{M}_{all}\}\leftarrow\mathop{param}$;
-20       $\mathop{Dist}_{max}\leftarrow\mathop{max}\{||\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||;i\in\{1,\cdots,m\}\}$;
-21       $\mathop{Dist}\leftarrow||\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$;
-22       if $||C_{\downarrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||<||C_{\uparrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$ then
+[…]
+[…]
+[…]
 22    $C_{dist}\leftarrow C_{\downarrow}$;
-24
+23
 24 else
 25    $C_{dist}\leftarrow C_{\uparrow}$;
-27
-28      while $\mathop{Dist}\geq\mathop{Dist}_{max}$& $||C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||\leq\mathop{Dist}$ do
+26
+27 while $\mathop{Dist}\geq\mathop{Dist}_{max}$ && $C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||\leq\mathop{Dist}$ do
 28    $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
-30             $\mathop{Dist}\leftarrow||\Delta\boldsymbol{g}_{t+1}^{(i)}-||$;
-31
+[…]
+30
 31 return $\Delta\boldsymbol{g}_{t+1}^{(i)}$;
-33
+32
 Algorithm 2 Similarity-based and distance-based adjustment functions in the Poison Pill Injection stage.
-Figure 4: Intuition behind distance-based adjustment in our augmentation method.
-Pill Adjustment.
-After the injection, we use a two-step adjustment to further adjust the pill, improving the generality against multiple detection metrics simultaneously.
-In this stage, we consider two prevailing detection metrics – distance and cosine similarity.
-To increase the cosine similarity between the poisoned model update and the benign model update in our method, we balance the magnitudes of both the poison pill’s parameters and the other benign parameters.
-Similarly, to minimize the distance discrepancy between the poisoned and benign model updates, we adjust the magnitude of the entire poisoned model update.
-Thus, we first use the similarity-based adjustment, then use the distance-based adjustment to balance the effectiveness and stealthiness poisoned model update.
-This two-step adjustment is particularly effective when combined with our method, which selectively poisons only a tiny subset of the model’s parameters.
-By altering just a few parameters, our method preserves a substantial number of benign parameters, which are crucial for making effective adjustments.
-As a result, the poisoned model update can bypass a wide range of defenses since they are typically designed based on the combination or variants of distance and cosine similarity metrics, and they usually do not anticipate such a focused and minimal interference in the model parameters.
-The details of the two seamless adjustments are as follows:
+The details of the two pill adjustment methods are presented as follows:
 Similarity-based Adjustment.
@@
 In the Distance-based Adjustment (Line 13--24 of Algorithm 2), we reuse the up-scaling factor $C_{\uparrow}$ and the down-scaling factor $C_{\downarrow}$ to adjust the magnitude of the entire poisoned model update.
-The intuition behind the Distance-based Adjustment is shown in Fig.
-4.
+The intuition behind the Distance-based Adjustment is shown in Figure 3.
 We first calculate the maximum distance between the normal model updates from the compromised clients and the estimated global model update in the current round.
@@
 We stop the scaling until the distance between the $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is smaller than $Dist_{max}$, or the scaling factor begins to increase such distance (reach the limit of the scaling).
-6 Evaluation
-This section assesses how our method enhances the effectiveness of existing FL poisoning attacks from multiple perspectives.
-We begin by evaluating the Augmentation Effectiveness of our method against four FL poisoning attacks, using nine prevailing defenses across three datasets, detailed in Section 6.2.
-Subsequently, we visualize the Stealthiness of our method under two prevailing detection metrics, as discussed in Section 6.3.
-Lastly, an Generality Analysis of our method is presented, which includes tests on various proportions of malicious clients, tests on both cross-silo and cross-device settings, and evaluates the impact of different pill search rules, outlined in Section 6.4.
-Our method significantly enhances the capabilities of existing FL poisoning attacks, successfully bypassing all $9$ baseline defenses in over 90% of cases, and increasing the error rates by up to seven times compared to the original attacks’ error rates.
-Moreover, it demonstrates robustness across varying data distributions, model architectures, proportions of malicious clients, and pill search rules.
-6.1 Evaluation Settings
-Attack, Defense, and Framework settings.
-In our experiments, we typically set the malicious proportion to $20\%$.
-We implement $9$ baseline aggregation rules, including FedAvg, FLTrust, Multi-Krum, Bulyan, Median, Trim, FLDetector, DnC, and Flame.
-We use our method to augment $4$ existing model poisoning attacks, including sign-flipping attack, Trim attack, Krum attack, and Min-Max attack.
-These attacks are chosen for their representativeness in illustrating the effectiveness of our method.
-We configure a 50-client FL system for both the MNIST and Fashion-MNIST datasets.
-For the CIFAR-10 dataset, a 30-client FL system is used.
-Our framework accommodates both cross-silo and cross-device settings.
-The entire framework is based on PyTorch [53].
+Appendix H Additional Experimental Configurations
 Model, Dataset, and Hyper-Parameters.
-In our experiments, we employ a four-layer Convolutional Neural Network (CNN) and a simplified version of AlexNet [54].
-The structures of the models and their corresponding pill blueprints are detailed in Appendix B.
-We evaluate our method on three widely-used datasets: MNIST [55], Fashion-MNIST [56], and CIFAR-10 [57].
-We use the CNN model on MNIST and Fashion-MNIST datasets, and the AlexNet [54] on CIFAR-10 dataset.
+In our experiments, we employ a four-layer Convolutional Neural Network (CNN) and a simplified version of AlexNet (alexnet).
+The structures of the models and their corresponding pill blueprints are detailed in Appendix D.
+We evaluate our method on three widely-used datasets: MNIST (lecun1998mnist), Fashion-MNIST (fashionmnist), and CIFAR-10 (cifar).
+We use the CNN model on MNIST and Fashion-MNIST datasets, and the AlexNet (alexnet) on CIFAR-10 dataset.
 Each experiment is repeated five times to ensure reliability, with the mean and standard deviation (std) of the results reported.
-TABLE III: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
-No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
-$\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003
-Sign-flipping Attack 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
-$\pm$0.023 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.002 $\pm$0.001 $\pm$0.003 $\pm$0.020 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.006 $\pm$0.003 $\pm$0.003
-0.667 0.115 0.764 0.379 0.523 0.314 0.646 0.543 0.122 0.754 0.430 0.522 0.311 0.688
-+ Poison Pill $\pm$0.089 $\pm$0.004 $\pm$0.049 $\pm$0.104 $\pm$0.091 $\pm$0.018 $\pm$0.061 $\pm$0.150 $\pm$0.006 $\pm$0.129 $\pm$0.057 $\pm$0.038 $\pm$0.038 $\pm$0.067
-Trim Attack 0.243 0.109 0.139 0.146 0.174 0.179 0.116 0.332 0.120 0.201 0.163 0.231 0.238 0.124
-$\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.006 $\pm$0.006 $\pm$0.003 $\pm$0.001 $\pm$0.022 $\pm$0.005 $\pm$0.018 $\pm$0.004 $\pm$0.008 $\pm$0.009 $\pm$0.003
-0.618 0.576 0.638 0.284 0.453 0.219 0.115 0.668 0.517 0.687 0.292 0.473 0.223 0.222
-+ Poison Pill $\pm$0.071 $\pm$0.057 $\pm$0.041 $\pm$0.040 $\pm$0.091 $\pm$0.010 $\pm$0.003 $\pm$0.033 $\pm$0.038 $\pm$0.036 $\pm$0.047 $\pm$0.047 $\pm$0.016 $\pm$0.128
-Krum Attack 0.116 0.109 0.189 0.201 0.172 0.137 0.786 0.128 0.116 0.235 0.276 0.217 0.160 0.947
-$\pm$0.002 $\pm$0.003 $\pm$0.022 $\pm$0.009 $\pm$0.008 $\pm$0.003 $\pm$0.087 $\pm$0.004 $\pm$0.003 $\pm$0.059 $\pm$0.003 $\pm$0.005 $\pm$0.003 $\pm$0.030
-0.735 0.155 0.715 0.422 0.578 0.310 0.637 0.716 0.151 0.737 0.468 0.730 0.334 0.690
-+ Poison Pill $\pm$0.032 $\pm$0.032 $\pm$0.132 $\pm$0.046 $\pm$0.057 $\pm$0.009 $\pm$0.074 $\pm$0.104 $\pm$0.004 $\pm$0.078 $\pm$0.017 $\pm$0.168 $\pm$0.031 $\pm$0.079
-Min-Max Attack 0.183 0.110 0.431 0.330 0.183 0.218 0.825 0.269 0.125 0.619 0.434 0.255 0.278 0.831
-$\pm$0.008 $\pm$0.002 $\pm$0.029 $\pm$0.015 $\pm$0.009 $\pm$0.009 $\pm$0.052 $\pm$0.026 $\pm$0.015 $\pm$0.050 $\pm$0.080 $\pm$0.012 $\pm$0.007 $\pm$0.049
-0.702 0.303 0.668 0.327 0.514 0.314 0.778 0.629 0.320 0.612 0.406 0.547 0.376 0.822
-+ Poison Pill $\pm$0.114 $\pm$0.201 $\pm$0.116 $\pm$0.074 $\pm$0.053 $\pm$0.047 $\pm$0.063 $\pm$0.114 $\pm$0.115 $\pm$0.040 $\pm$0.065 $\pm$0.072 $\pm$0.119 $\pm$0.036
 IID and Non-IID Data Settings.
@@
 For IID data setting, we uniformly split all the training data into $K$ shards, and distribute each shard to a random client.
-For non-IID data setting, we utilize the non-IID degree $p$ as defined in prior studies [4, 30].
+For non-IID data setting, we utilize the non-IID degree $p$ as defined in prior studies (fang2020local; guo2021siren).
 A higher $p$ indicates greater data heterogeneity among the clients.
@@
 This approach ensures that there is no overlap between server’s data and client’s data.
-Configurations of Dynamic Patterns in Our Method.
+Configurations of Dynamic Patterns.
 As outlined in Section 4.2, we design six dynamic patterns for the pill search.
@@
 Given that the model poisoning attacks discussed are all untargeted, higher error rates indicate more effective attacks.
-To assess the stealthiness of our method in delivering malicious updates, we employ two metrics: 1) cosine similarity score: measures alignment with the server’s model update in FLTrust; 2) distance score: used in Multi-Krum to evaluate the closeness of poisoned updates to benign updates.
-TABLE IV: Error rates under cross-silo setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
-No Attack 0.106 0.104 0.103 0.108 0.127 0.107 0.116 0.111 0.119 0.113 0.113 0.140 0.114 0.123
-$\pm$ 0.003 $\pm$0.003 $\pm$0.003 $\pm$0.004 $\pm$0.001 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.001 $\pm$0.002 $\pm$0.005 $\pm$0.002 $\pm$0.004
-Sign-flipping Attack 0.964 0.109 0.108 0.110 0.130 0.108 0.117 0.909 0.119 0.114 0.119 0.144 0.120 0.125
-$\pm$0.017 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.001 $\pm$0.005 $\pm$0.045 $\pm$0.002 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.002
-0.320 0.116 0.162 0.151 0.323 0.148 0.699 0.269 0.120 0.239 0.164 0.364 0.168 0.242
-+ Poison Pill $\pm$0.080 $\pm$0.007 $\pm$0.027 $\pm$0.010 $\pm$0.029 $\pm$0.007 $\pm$0.082 $\pm$0.174 $\pm$0.003 $\pm$0.101 $\pm$0.013 $\pm$0.031 $\pm$0.012 $\pm$0.198
-Trim Attack 0.112 0.111 0.111 0.115 0.132 0.114 0.116 0.125 0.115 0.121 0.125 0.153 0.122 0.122
-$\pm$0.002 $\pm$0.005 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.006 $\pm$0.001 $\pm$0.004 $\pm$0.005 $\pm$0.002 $\pm$0.002
-0.508 0.139 0.334 0.126 0.284 0.127 0.120 0.528 0.148 0.455 0.143 0.287 0.146 0.136
-+ Poison Pill $\pm$0.128 $\pm$0.012 $\pm$0.120 $\pm$0.006 $\pm$0.040 $\pm$0.004 $\pm$0.003 $\pm$0.051 $\pm$0.018 $\pm$0.151 $\pm$0.003 $\pm$0.023 $\pm$0.004 $\pm$0.012
-Krum Attack 0.107 0.108 0.114 0.123 0.141 0.112 0.668 0.116 0.117 0.124 0.138 0.173 0.122 0.410
-$\pm$0.004 $\pm$0.005 $\pm$0.001 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.134 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.001 $\pm$0.005 $\pm$0.003 $\pm$0.352
-0.183 0.118 0.283 0.161 0.362 0.146 0.631 0.428 0.127 0.280 0.187 0.415 0.182 0.704
-+ Poison Pill $\pm$0.039 $\pm$0.003 $\pm$0.210 $\pm$0.015 $\pm$0.027 $\pm$0.008 $\pm$0.089 $\pm$0.190 $\pm$0.005 $\pm$0.064 $\pm$0.012 $\pm$0.057 $\pm$0.009 $\pm$0.091
-Min-Max Attack 0.117 0.108 0.118 0.135 0.142 0.128 0.111 0.124 0.119 0.142 0.166 0.162 0.145 0.136
-$\pm$0.004 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.009 $\pm$0.008 $\pm$0.004 $\pm$0.003 $\pm$0.012 $\pm$0.003 $\pm$0.007 $\pm$0.004 $\pm$0.004 $\pm$0.002
-0.439 0.129 0.361 0.136 0.343 0.150 0.715 0.521 0.136 0.339 0.153 0.368 0.184 0.335
-+ Poison Pill $\pm$0.140 $\pm$0.009 $\pm$0.245 $\pm$0.015 $\pm$0.032 $\pm$0.006 $\pm$0.096 $\pm$0.073 $\pm$0.011 $\pm$0.202 $\pm$0.009 $\pm$0.048 $\pm$0.017 $\pm$0.185
-Figure 5: Comparison of Multi-Krum distance score between benign updates and malicious updates when using original poisoning attacks with and without our method.
-TABLE V: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on CIFAR-10 dataset.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame
-No Attack 0.488 0.480 0.507 0.469 0.551 0.456 0.445 0.494 0.491 0.486 0.474 0.499 0.498 0.581 0.502 0.463 0.506 0.532
-Sign-flipping Attack 0.898 0.479 0.580 0.539 0.621 0.461 0.468 0.497 0.509 0.905 0.514 0.511 0.622 0.658 0.573 0.502 0.603 0.533
-+ Poison Pill 0.739 0.880 0.929 0.694 0.707 0.699 0.536 0.899 0.706 0.879 0.861 0.898 0.677 0.766 0.688 0.566 0.900 0.675
-Trim Attack 0.482 0.509 0.489 0.536 0.623 0.514 0.456 0.459 0.501 0.571 0.493 0.608 0.595 0.653 0.549 0.481 0.482 0.506
-+ Poison Pill 0.853 0.877 0.883 0.654 0.674 0.662 0.513 0.899 0.542 0.890 0.862 0.906 0.772 0.688 0.639 0.518 0.893 0.621
-Krum Attack 0.473 0.541 0.471 0.568 0.540 0.510 0.455 0.802 0.500 0.485 0.506 0.497 0.522 0.647 0.519 0.481 0.899 0.501
-+ Poison Pill 0.701 0.896 0.900 0.765 0.756 0.643 0.529 0.890 0.872 0.724 0.849 0.900 0.675 0.748 0.647 0.580 0.885 0.873
-Min-Max Attack 0.450 0.504 0.469 0.507 0.579 0.465 0.514 0.525 0.525 0.478 0.502 0.493 0.568 0.636 0.603 0.478 0.482 0.488
-+ Poison Pill 0.752 0.712 0.902 0.775 0.802 0.640 0.545 0.902 0.811 0.661 0.646 0.886 0.674 0.783 0.661 0.527 0.907 0.799
-Figure 6: Comparison of cosine similarity scores between original attack with and without our method under FLTrust.
-Figure 7: Comparison of the cosine similarity scores with the server model of the original attack and our different augmentation patterns in the entire training period under FLTrust.
-Figure 8: Comparison of error rates between original poisoning attacks and our method with two different pill search methods.
-6.2 Augmentation Effectiveness
-In this section, we present a comprehensive analysis of our method’s augmentation effectiveness on Fashion-MNIST dataset within a 50-client cross-silo FL system, in which 20% clients are malicious.
-We evaluate our method on both IID data and non-IID data.
-Our method successfully augments all the baseline attacks with more than $0.25$ average error rate increase, showing our method’s effectiveness and high compatibility.
-Results on IID Data.
-The error rates of four baseline FL poisoning attacks, with and without our method, are shown in the left half of Table III.
-Our method enhances the error rates of the existing poisoning attacks in $23$ out of $28$ scenarios, against FedAvg and five defenses.
-The maximum increase in error rate is $0.658$, and the average increase reaches $0.274$.
-This substantial elevation from the attack-free baseline error rate of $0.109$ underscores our method’s capability to significantly compromise existing defenses’ integrity.
-Following are individual improvements of our method on different baseline attacks:
+To assess the stealthiness of our method in delivering malicious updates, we employ two metrics: 1) cosine similarity score, measuring alignment with the server’s model update in FLTrust; 2) distance score, used in Multi-Krum to evaluate the closeness of poisoned updates to benign updates.
+Appendix I Additional Augmentation Performance Analysis
+Following are individual improvements of our method on different baseline attacks in IID data setting:
   • Sign-flipping attack: Its original version achieves a high error rate due to its aggressive and brute design, but it is effective only under FedAvg.
@@
   • Min-Max attack: With our method, the Min-Max Attack shows a comprehensive improvement against all defenses except for a slight decrease against Bulyan, achieving an overall average error rate increase of $0.222$.
-Results on Non-IID Data.
-Evaluations on non-IID data further validate the effectiveness of our method, demonstrating its superiority in $23$ out of $28$ cases.
-The highest error rate increase reaches $0.637$, with an average increase of $0.281$.
-Although there is a slight reduction in maximal error rate increase in the non-IID data setting, these results still demonstrate our method’s ability to effectively enhance attacks in more complex and heterogeneous data environments.
-The detailed improvements for a specific attack are shown as follows:
+Similarly, the detailed improvements for a specific attack in the non-IID data setting shown as follows:
   • Sign-flipping attack: Our method helps the sign-flipping attack achieve an average error rate increase of $0.404$, which is similar to the error rate increase on IID data.
@@
 Although this error rate increase is lower than that in the IID data setting, it remains higher than the error rate increase caused by its original version.
-All attacks augmented by our method can bypass all baseline defenses, including FLTrust and FLDetector, with the exception of the sign-flipping attack.
-Notably, the Min-Max attack demonstrates superior effectiveness in non-IID data settings, achieving significant improvements compared to its performance on IID data.
-Other attacks also exhibit similar error rate improvements relative to their results on IID data, indicating that our method maintains its robustness and effectiveness in more complex data environments.
-6.3 Stealthiness Analysis
-To further analyze the performance of our method, we analyze its stealthiness during the training process of the FL system, focusing on how our method influences the distance scores and cosine similarity scores of exisitng FL poisoning attacks.
-The results indicate that our method can make malicious clients appear as benign or even more “benign” than genuine benign clients.
-This significant increase in stealthiness is a result of the pill design with distance-based adjustment and similarity-based adjustment techniques in our method.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.028 0.051 0.029 0.029 0.045 0.029 0.025 0.029 0.042 0.030 0.029 0.041 0.029 0.022
+Sign-Flipping 0.934 0.059 0.038 0.055 0.055 0.036 0.025 0.886 0.073 0.041 0.052 0.059 0.041 0.026
+[…]
+Trim Attack 0.257 0.065 0.182 0.103 0.106 0.123 0.022 0.418 0.059 0.295 0.209 0.245 0.310 0.021
+[…]
+Krum Attack 0.033 0.061 0.067 0.154 0.188 0.043 0.759 0.034 0.058 0.130 0.297 0.191 0.052 0.908
+[…]
+Min-Max Attack 0.307 0.082 0.693 0.731 0.341 0.255 0.915 0.359 0.161 0.718 0.993 0.381 0.320 0.853
+[…]
+Table 8: Error rates on MNIST under cross-silo setting with 20% malicious clients in the 50-client FL system.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
+No Attack 0.107 0.111 0.108 0.105 0.138 0.106 0.113 0.124 0.115 0.118 0.164 0.116
+Sign-Flipping 0.940 0.116 0.110 0.128 0.165 0.121 0.905 0.124 0.118 0.136 0.184 0.134
+[…]
+Trim Attack 0.240 0.110 0.151 0.148 0.207 0.178 0.340 0.120 0.228 0.190 0.237 0.245
+[…]
+Krum Attack 0.117 0.112 0.172 0.238 0.169 0.132 0.126 0.121 0.204 0.296 0.222 0.158
+[…]
+Min-Max Attack 0.146 0.111 0.382 0.324 0.183 0.185 0.191 0.147 0.621 0.426 0.245 0.279
+[…]
+Table 9: Error rates on Fashion-MNIST under cross-device setting with 20% malicious clients in 50-client FL system.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
+No Attack 0.110 0.106 0.107 0.108 0.139 0.106 0.115 0.117 0.115 0.117 0.164 0.112
+Sign-Flipping 0.929 0.111 0.108 0.111 0.153 0.117 0.902 0.118 0.115 0.120 0.175 0.134
+[…]
+Trim Attack 0.112 0.114 0.111 0.118 0.153 0.113 0.129 0.125 0.128 0.129 0.185 0.122
+[…]
+Krum Attack 0.110 0.113 0.115 0.128 0.144 0.113 0.121 0.116 0.123 0.135 0.183 0.120
+[…]
+Min-Max Attack 0.116 0.111 0.116 0.127 0.145 0.122 0.121 0.116 0.123 0.135 0.183 0.120
+[…]
+Table 10: Error rates on Fashion-MNIST under cross-device setting with 10% malicious clients in 50-client FL system.
+Appendix J Additional Stealthiness Analysis
+Figure 6: Comparison of cosine similarity scores between original attack with and without our method under FLTrust.
+Figure 7: Comparison of the cosine similarity scores with the server model of the original attack and our different augmentation patterns in the entire training period under FLTrust.
 Distance Score Analysis.
@@
 In contrast, original attacks like the Trim and sign-flipping attacks display distance scores that were significantly higher or lower than those of benign updates, indicating either detected by Multi-Krum (higher scores) or underutilized attack capacities (lower scores).
-Our method also has lower distance score variance in the early FL training period, representing that our method provides more steady attack efficacy in the FL’s critical training period [58, 59] by fully utilizing the attack capacities while being undetected.
-More details are shown in Appendix D.
+Our method also has a lower distance score variance in the early FL training period, representing that our method provides more steady attack efficacy in the FL’s critical training period (yan2023defl; yan2022seizing) by fully utilizing the attack capacities while being undetected.
+Additionally, our method also achieves two more improvements.
+First, our method causes the global model to degrade earlier compared to the original attacks, further demonstrating the effectiveness of our augmentation.
+Second, our method significantly increases the discrepancy between benign client updates as the communication rounds increase.
+While original attacks can bypass detection in some cases, the discrepancy between benign client updates remains steady, illustrating the lower impact of malicious clients.
+In contrast, our method consistently increases the discrepancy among benign clients, highlighting its penetrating effectiveness in its influence on benign clients’ local training.
 Cosine Similarity Score Analysis.
@@
 This consistency not only ensures the successful insertion of the pill in any specific round but also secures pill’s long-lasting presence in the global model.
-6.4 Generality Analysis
-In this section, we further discuss the generality of our method from four perspectives: malicious client proportion, client participation frequency, datasets & model architectures, and pill search algorithm.
-The results indicate that our method maintains its augmentation effectiveness consistently, even as these conditions change, demonstrating its reliability and wide applicability in augmenting FL poisoning attacks.
-Impact of The Malicious Client’s Proportion.
-We first assess the effectiveness of our method in both IID and non-IID cross-silo FL system with only $10$% of clients compromised, as shown in Table IV.
-This setup revealed that all baseline model poisoning attacks yield lower error rates on the global model compared with scenarios with $20$% compromised clients.
-While the increase in error rates was less than those in the $20$% compromised client scenario, our method still effectively raises the global model’s error rates in $25$/$26$ out of $28$ cases (IID/non-IID setting).
-The maximum increase in error rates reached $0.403$, with an average increase of $0.144$.
-This average is notably higher (>$2$x higher) than the error rates observed in attack-free FL conditions.
-Specifically, our method help sign-flipping/Trim/Krum/Min-Max attacks achieve an average error rate increase of $0.133$/$0.094$/$0.136$/$0.209$.
-More detailed results are presented in Appendix G.
-Impact of The Client Participation Frequency.
-We then extend the evaluation of our method to a cross-device FL system, where only $40$% of clients are selected for participation in each communication round.
-This setup results in less frequent participation from each client and a fluctuating proportion of malicious clients across different rounds.
-The maximum error rate increase with our method is $0.639$, with an average increase across different attacks and defenses of $0.279$.
-These results are consistent with those from the cross-silo FL system, underscoring our method’s effectiveness and generality across different FL configurations.
-This evaluation demonstrates our method’s robust performance and adaptability, not only in a controlled cross-silo environment but also under the more various conditions in cross-device FL systems.
-More details are presented in Appendix F.
-Impact of The Datasets and Model Architectures.
-Following the evaluation with the Fashion-MNIST dataset, we test our method on the MNIST and CIFAR-10 dataset, employing the four-layer CNN model and the AlexNet model to further verify our method’s generality across different datasets.
-The collective results show that our method performs even better with larger datasets or more complex machine learning models.
-This trend confirms the generality of our method by revealing its capability to maintain consistent performance enhancements regardless of the dataset or model complexity involved.
-Specifically, our method help all the four baseline attacks bypass all the nine baselines on CIFAR-10 dataset, achieving $0.288$ error rate increase on average, presented in Table V.
-More detailed results on MNIST dataset are shown in Appendix E.
-Impact of The Pill Search Algorithm.
-We conduct a final evaluation to assess the importance and effectiveness of the "approximate max pill search" algorithm used in our method.
-This is contrasted against a newly devised "approximate min pill search" algorithm, which targets the least important parameters within the target model.
-Figure 8 illustrates the error rates achieved by the "approximate max pill search", the "approximate min pill search", and the original model poisoning attacks.
-The "approximate max pill search" algorithm outperforms the "approximate min pill search" in $41$ out of $56$ cases (approximately $73$%), underscoring its effectiveness in leveraging the most influential parameters to enhance attack impacts.
-Despite its lower efficacy, the "approximate min pill search" still manage to surpass the original attacks in $41$ out of $56$ cases (approximately $73$%).
-This demonstrates the generality of our method across different pill search algorithms.
-7 Discussion
-7.1 Our Method against Possible Adaptive Defense
-To further evaluate the robustness of our method when defenses are aware of the attack strategies (white-box scenario), we develop an adaptive defense named DSTrust, which enhances the FLTrust’s mechanism.
-DSTrust incorporates both distance and cosine similarity scores into a unified trust score calculation, directly countering our method’s two-step adjustment approach.
-The round-$t$ trust score of client $i$ in DSTrust is calculated as follows:
-$TS_{i}=ReLU(\frac{\mathop{cos}(\Delta\boldsymbol{g}_{t}^{(i)},\Delta\boldsymbol{g}_{t}^{s})}{||\Delta\boldsymbol{g}_{t}^{(i)}-\Delta\boldsymbol{g}_{t}^{s}||}),$ (3)
-where $\Delta\boldsymbol{g}_{t}^{(i)}$ represents the model update from client $i$ and $\Delta\boldsymbol{g}_{t}^{s})$ represents server’s model update.
-By integrating both cosine similarity and distance metrics, DSTrust provides a more comprehensive defense approach compared with FLTrust.
-This dual consideration allows DSTrust to effectively mitigate attacks that manipulate either of these metrics to bypass defenses.
-TABLE VI: Error rates under cross-silo setting against the new adaptive defense – DSTrust – with and without our method on Fashion-MNIST dataset (20% malicious clients).
-Data Distribution IID Non-IID
-[…]
-[…]
-[…]
-[…]
- Poison Pill
-No Attack 0.108 0.116
-[…]
- Attack 0.111 0.129 0.110 0.131
-Trim Attack 0.109 0.629 0.115 0.630
-Krum Attack 0.111 0.140 0.120 0.128
-Min-Max Attack 0.127 0.167 0.143 0.327
-Table VI details the error rates for four baseline FL poisoning attacks both with and without our method against the DSTrust defense on the Fashion-MNIST dataset within a 50-client FL system, where 20% of clients are malicious.
-These tests were conducted under both IID and non-IID data environments.
-DSTrust effectively neutralizes the four baseline poisoning attacks when our method is not applied, highlighting its robustness as a defense mechanism.
-Despite DSTrust’s integration of both cosine similarity and distance metrics in its defense strategy, it fails to counteract the augmented attacks when our method is employed.
-Notably, our method achieves a maximum error rate increase of $0.521$, and an average error rate increase of $0.173$ across all $8$ test scenarios.
-These results demonstrate that merely understanding the adjustment strategies of our method, and subsequently integrating corresponding defense metrics, does not fundamentally negate the effectiveness of our method.
-Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics into DSTrust, it remains insufficient to thwart the enhanced capabilities of our method.
-7.2 Limitations and Future Work
-Our method significantly enhances non-state-of-the-art (non-SOTA) model poisoning attacks, enabling them to SOTA results against various prevalent defenses.
-This is accomplished through a pill-based, attack-agnostic augmentation pipeline.
-We not only demonstrate our method’s capabilities but also expose fundamental vulnerabilities within the current designs of defense mechanisms.
-For future attacks in FL, it is essential for attackers to meticulously evaluate the importance of each parameter in their implementation.
-By targeting specific subsets of parameters, attackers can devise more flexible and adaptive attacks, improving stealthiness and complicating defense efforts.
-As for future defenses, while individually checking each parameter might seem viable, its practical deployment is hindered by high overheads, making it infeasible in real-world applications.
-Thus, there is a pressing need for more sophisticated defenses that can conduct fine-grained analyses of the roles of different parameters in neural networks, while executing without imposing prohibitive computational costs.
-8 Conclusion
-In this paper, we propose a novel attack-agnostic augmentation method to enhance existing poisoning attacks in FL by concentrating attacks into a pill (a tiny subnet).
-Our approach is constructed with three stages, including pill construction, pill poisoning, and pill injection.
-Accordingly, we first use a dynamic pill search algorithm to identify the pill blueprint.
-Then we poison the pill using existing FL poisoning attacks, and carefully inject the poison pill into the target model with two pill-related masks and a two-step adjustment.
-Our method enables existing poisoning attacks to achieve more than $2$x the error rates on average compared with their original implementations.
-The effectiveness of our method in exploiting and exacerbating the inherent weaknesses of current FL defenses highlights the critical need for more refined detection measures for FL.
-Appendix A Additional Details of The Baseline Defenses
-Krum and Multi-Krum (MKrum) [20].
-Krum uses a distance score as the metric.
-In each round, the Krum server sums the distances between each client update $\boldsymbol{g}_{t}^{(i)}$ and its $K-m-2$ neighbors, and uses these sums as the scores for all the clients.
-The Krum server then selects the client’s model update with the lowest score.
-Multi-Krum is a variant of Krum that uses iterative Krum to pick multiple candidates for aggregation.
-Coordinate-wise Median (Median) [26].
-Coordinate-wise Median (Median) uses the per-parameter median values of the model updates from the clients as the aggregated global model update, which is then used to generate the next-round global model.
-Trimmed Mean (Trim) [26]. Trimmed Mean (Trim) calculates per-parameter trimmed mean values of the client model updates and packs them as the global model update.
-Bulyan [27].
-Bulyan is a combination of Krum and Trim.
-It first uses the Krum-based method to select multiple candidates, and uses the per-parameter trimmed mean values of the candidate model updates as the final global model update.
-FLTrust [21].
-FLTrust trains a server model with a small root dataset.
-In each round, it computes the clipped cosine similarities between the server model update and client updates as trust scores, and then uses the trust scores as weights to aggregate all the normalized client model updates.
-FLDetector (FLD) [39].
-FLDetector filters out malicious clients by checking the multi-round consistency of all client updates.
-Malicious updates typically have lower consistency compared to benign ones.
-Flame [23]. Flame utilizes HDBSCAN-based[60] dynamic clustering to filtered out malicious clients, and aggregates median-clipped benign updates with adaptive noise as the global model update.
-Appendix B Additional Details of Concrete Pill Blueprints
-Table VII and Table VIII illustrate the model structures of the CNN model and the simplified AlexNet, with their corresponding pill’s blueprints.
-TABLE VII: Architectures of the original CNN model and the corresponding pill blueprint.
-[…]
-[…]
-Pill Blueprint
-Input $28\times 28\times 1$ $28\times 28\times 1$
-Conv2d $3\times 3\times 30$ $3\times 3\times 1$
-ReLU -1 -
-MaxPool2d $2\times 2$ $2\times 2$
-Conv2d $3\times 3\times 50$ $3\times 3\times 1$
-ReLU - -
-MaxPool2d $2\times 2$ $2\times 2$
-Linear $1250\times 100$ $25\times 10$
-ReLU - -
-Linaer $100\times 10$ $10$
-[…]
-  • 1
-    “-” represents that the model has this layer with no specified configuration.
-  • 2
-    $\times$” represents that the model does not contain this layer.
-TABLE VIII: Example architectures of original AlexNet and the corresponding pill blueprint.
-[…]
-[…]
-Pill Blueprint
-Input $32\times 32\times 3$ $32\times 32\times 3$
-Conv2d $11\times 11\times 48$ $11\times 11\times 1$
-ReLU - -
-MaxPool2d $3\times 3$ $3\times 3$
-Conv2d $3\times 3\times 96$ $3\times 3\times 1$
-ReLU - -
-MaxPool2d $3\times 3$ $3\times 3$
-Conv2d $3\times 3\times 192$ $3\times 3\times 1$
-ReLU - -
-Conv2d $3\times 3\times 192$ $3\times 3\times 1$
-ReLU - -
-Conv2d $3\times 3\times 128$ $3\times 3\times 1$
-ReLU - -
-MaxPool2d $3\times 3$ $3\times 3$
-Linear $4608\times 1024$ $36\times 1$
-ReLU - -
-Linear $1024\times 512$ $1\times 10$
-ReLU - -
-Linear $512\times 10$ $10$
-Softmax - $\times$
-Appendix C Additional Details of The Dynamic Patterns in Our Method
-We first design three searching strategies, including one-time searching strategy, repeated searching strategy, and adaptive searching strategy.
-In the one-time searching strategy, we search the pill based on the initial global model using the “approximate max pill search” algorithm introduced in §5.2, and keep this pill unchanged in the whole FL training.
-The one-time searching strategy benefits the formation of pill in the global model, while the initial pill may be less effective with the increasing training rounds of the global model, due to the changing importance of the model parameters.
-On the contrary, the repeated searching strategy runs the ‘approximate max pill search” algorithm in every training round.
-The repeated searching strategy can help our method modify more parameters in the global model, and make the pill less traceable.
-While the attacking effects may be reduced due to the constantly changing pill.
-Considering advantages and disadvantages of both the one-time searching strategy and the repeated searching strategy, we design a more flexible searching strategy, termed as “adaptive searching strategy”.
-In adaptive searching strategy, our method searches the new pill only when the pill is not successfully injected into the global model in the last round.
-The condition:
-[…]
-should be satisfied to trigger the new subnetwork searching on malicious client $i$, where $C_{search}$ is set as $0.94$ in the experiments.
-The adaptive searching strategy is a mroe moderate version of repeated searching.
-Since the three searching strategies have their unique advantages, we investigate different combinations of them in the experiments.
-We further divide the neural network into Feature Extractor (FE) and Classifier (CLS).
-Refer to the CNN model we used, the convolutional layers are regarded as FE, and the linear layers are regarded as CLS.
-We use different searching strategies in FE and CLS, respectively.
-In all the nine combinations, we test and keep six of them, noted as Pattern1 to Pattern6, shown in Table II in § 5.2.
-Such six patterns construct the entire dynamic pattern set used in our method.
-TABLE IX: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on MNIST dataset.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
-No Attack 0.028 0.051 0.029 0.029 0.045 0.029 0.025 0.029 0.042 0.030 0.029 0.041 0.029 0.022
-Sign-flipping Attack 0.934 0.059 0.038 0.055 0.055 0.036 0.025 0.886 0.073 0.041 0.052 0.059 0.041 0.026
-+ Poison Pill 0.353 0.093 0.454 0.283 0.268 0.173 0.588 0.431 0.059 0.605 0.349 0.333 0.217 0.713
-Trim Attack 0.257 0.065 0.182 0.103 0.106 0.123 0.022 0.418 0.059 0.295 0.209 0.245 0.310 0.021
-+ Poison Pill 0.416 0.109 0.469 0.252 0.247 0.117 0.026 0.581 0.065 0.672 0.358 0.324 0.092 0.051
-Krum Attack 0.033 0.061 0.067 0.154 0.188 0.043 0.759 0.034 0.058 0.130 0.297 0.191 0.052 0.908
-+ Poison Pill 0.326 0.082 0.585 0.266 0.272 0.169 0.632 0.528 0.062 0.556 0.350 0.321 0.210 0.746
-Min-Max Attack 0.307 0.082 0.693 0.731 0.341 0.255 0.915 0.359 0.161 0.718 0.993 0.381 0.320 0.853
-+ Poison Pill 0.402 0.106 0.518 0.273 0.262 0.218 0.766 0.534 0.077 0.707 0.369 0.318 0.194 0.861
-TABLE X: Error rates under cross-device setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
-No Attack 0.107 0.111 0.108 0.105 0.138 0.106 0.113 0.124 0.115 0.118 0.164 0.116
-$\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.008 $\pm$0.006 $\pm$0.002 $\pm$0.009 $\pm$0.005
-Sign-flipping Attack 0.940 0.116 0.110 0.128 0.165 0.121 0.905 0.124 0.118 0.136 0.184 0.134
-$\pm$0.026 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.007 $\pm$0.004 $\pm$0.031 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.007 $\pm$0.006
-0.591 0.117 0.749 0.357 0.589 0.225 0.573 0.125 0.665 0.379 0.662 0.277
-+ Poison Pill $\pm$0.177 $\pm$0.004 $\pm$0.076 $\pm$0.057 $\pm$0.048 $\pm$0.026 $\pm$0.140 $\pm$0.004 $\pm$0.111 $\pm$0.018 $\pm$0.131 $\pm$0.012
-Trim Attack 0.240 0.110 0.151 0.148 0.207 0.178 0.340 0.120 0.228 0.190 0.237 0.245
-$\pm$0.018 $\pm$0.004 $\pm$0.010 $\pm$0.002 $\pm$0.014 $\pm$0.004 $\pm$0.048 $\pm$0.002 $\pm$0.025 $\pm$0.016 $\pm$0.016 $\pm$0.011
-0.620 0.492 0.620 0.228 0.424 0.232 0.654 0.533 0.679 0.324 0.483 0.226
-+ Poison Pill $\pm$0.051 $\pm$0.023 $\pm$0.025 $\pm$0.025 $\pm$0.042 $\pm$0.035 $\pm$0.037 $\pm$0.041 $\pm$0.049 $\pm$0.098 $\pm$0.098 $\pm$0.025
-Krum Attack 0.117 0.112 0.172 0.238 0.169 0.132 0.126 0.121 0.204 0.296 0.222 0.158
-$\pm$0.002 $\pm$0.004 $\pm$0.010 $\pm$0.005 $\pm$0.009 $\pm$0.003 $\pm$0.005 $\pm$0.004 $\pm$0.031 $\pm$0.011 $\pm$0.014 $\pm$0.006
-0.681 0.138 0.740 0.362 0.572 0.258 0.604 0.141 0.750 0.372 0.649 0.277
-+ Poison Pill $\pm$0.057 $\pm$0.015 $\pm$0.092 $\pm$0.073 $\pm$0.167 $\pm$0.018 $\pm$0.125 $\pm$0.009 $\pm$0.081 $\pm$0.035 $\pm$0.184 $\pm$0.013
-Min-Max Attack 0.146 0.111 0.382 0.324 0.183 0.185 0.191 0.147 0.621 0.426 0.245 0.279
-$\pm$0.005 $\pm$0.002 $\pm$0.036 $\pm$0.012 $\pm$0.011 $\pm$0.006 $\pm$0.014 $\pm$0.024 $\pm$0.112 $\pm$0.072 $\pm$0.013 $\pm$0.007
-0.651 0.244 0.718 0.312 0.503 0.249 0.670 0.229 0.621 0.349 0.581 0.386
-+ Poison Pill $\pm$0.082 $\pm$0.104 $\pm$0.059 $\pm$0.026 $\pm$0.060 $\pm$0.014 $\pm$0.123 $\pm$0.098 $\pm$0.030 $\pm$0.047 $\pm$0.161 $\pm$0.141
-TABLE XI: Error rates under cross-device setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
-Data Distribution IID Non-IID
-Attack FedAvg FLTrust MKrum Bulyan Median Trim FedAvg FLTrust MKrum Bulyan Median Trim
-No Attack 0.110 0.106 0.107 0.108 0.139 0.106 0.115 0.117 0.115 0.117 0.164 0.112
-$\pm$0.004 $\pm$0.004 $\pm$0.005 $\pm$0.003 $\pm$0.008 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.002
-Sign-flipping Attack 0.929 0.111 0.108 0.111 0.153 0.117 0.902 0.118 0.115 0.120 0.175 0.134
-$\pm$0.026 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.025 $\pm$0.004 $\pm$0.034 $\pm$0.001 $\pm$0.004 $\pm$0.003 $\pm$0.008 $\pm$0.008
-0.195 0.114 0.170 0.138 0.347 0.137 0.330 0.124 0.165 0.148 0.483 0.161
-+ Poison Pill $\pm$0.032 $\pm$0.003 $\pm$0.071 $\pm$0.005 $\pm$0.059 $\pm$0.004 $\pm$0.135 $\pm$0.006 $\pm$0.016 $\pm$0.007 $\pm$0.225 $\pm$0.009
-Trim Attack 0.112 0.114 0.111 0.118 0.153 0.113 0.129 0.125 0.128 0.129 0.185 0.122
-$\pm$0.003 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.021 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.011 $\pm$0.003
-0.369 0.138 0.212 0.128 0.310 0.140 0.589 0.154 0.300 0.139 0.351 0.156
-+ Poison Pill $\pm$0.147 $\pm$0.015 $\pm$0.066 $\pm$0.005 $\pm$0.038 $\pm$0.014 $\pm$0.046 $\pm$0.017 $\pm$0.100 $\pm$0.006 $\pm$0.056 $\pm$0.015
-Krum Attack 0.110 0.113 0.115 0.128 0.144 0.113 0.121 0.116 0.123 0.135 0.183 0.120
-$\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
-0.164 0.117 0.157 0.143 0.371 0.142 0.229 0.126 0.249 0.146 0.374 0.157
-+ Poison Pill $\pm$0.038 $\pm$0.003 $\pm$0.044 $\pm$0.010 $\pm$0.034 $\pm$0.005 $\pm$0.069 $\pm$0.002 $\pm$0.167 $\pm$0.004 $\pm$0.023 $\pm$0.005
-Min-Max Attack 0.116 0.111 0.116 0.127 0.145 0.122 0.121 0.116 0.123 0.135 0.183 0.120
-$\pm$0.002 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.006 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
-0.351 0.124 0.299 0.135 0.343 0.146 0.342 0.138 0.292 0.148 0.417 0.166
-+ Poison Pill $\pm$0.204 $\pm$0.019 $\pm$0.110 $\pm$0.004 $\pm$0.070 $\pm$0.019 $\pm$0.076 $\pm$0.018 $\pm$0.087 $\pm$0.009 $\pm$0.050 $\pm$0.010
-Appendix D Additional Stealthiness Analysis
-As shown in Figure 5, in addition to the minimal difference between the benign updates and the augmented malicious updates, our method also achieves two more improvements.
-First, our method causes the global model to degrade earlier compared to the original attacks, further demonstrating the effectiveness of our augmentation.
-Second, our method significantly increases the discrepancy between benign client updates as the communication rounds increase.
-While original attacks can bypass detection in some cases, the discrepancy between benign client updates remains steady, illustrating the lower impact of malicious clients.
-In contrast, our method consistently increases the discrepancy among benign clients, highlighting its penetrating effectiveness in its influence on benign clients’ local training.
-Appendix E Additonal Details on MNIST and CIFAR-10 Dataset
-The detailed results on MNIST and CIFAR-10 dataset are presented in Table IX (MNIST dataset) and Table V (CIFAR-10 dataset), respectively.
+Appendix K Additional Details on MNIST and CIFAR-10 Dataset
+The detailed results on MNIST and CIFAR-10 datasets are presented in Table 8 (MNIST dataset) and Table 3 (CIFAR-10 dataset), respectively.
 For MNIST dataset, the highest error rate increase achieved using our method is $0.518$, with an average increase of $0.121$.
@@
 Despite the reduced average error rate increase, it remains significant, especially considering the MNIST dataset’s lower baseline error rates (below 0.070).
-On CIFAR-10 dataset, our method help existing FL poisoning attacks outperform their original versions in $71$ of the $72$ scenarios, with an average error rate increase of over $0.288$.
+On CIFAR-10 dataset, our method helps existing FL poisoning attacks outperform their original versions in $71$ of the $72$ scenarios, with an average error rate increase of over $0.288$.
 Specifically, our method facilitates at least a $0.212$ increase in error rates against FLTrust, outperforming the results in the same settings on the Fashion-MNIST dataset.
-Appendix F Additional Results in Cross-device FL System
+Appendix L Additional Results in Cross-device FL System
 After evaluating our method in the 50-client cross-silo FL system, we further test it in the 50-client cross-device FL system.
-Table X presents the error rates under the cross-device FL setting using the “approximate max pill search” algorithm on both IID and non-IID data.
+Table 9 presents the error rates under the cross-device FL setting using the “approximate max pill search” algorithm on both IID and non-IID data.
 We report the highest error rates among the results of six dynamic patterns, with the malicious client proportion set to $20\%$.
@@
 For the Min-Max attack, the highest error rate increase reaches $0.505$, with an average increase of $0.272$.
-These improvements are consistent with the error rates observed under the cross-silo FL setting using the "approximate max pill search" algorithm on IID data.
+These improvements are consistent with the error rates observed under the cross-silo FL setting using the ”approximate max pill search” algorithm on IID data.
 Results on non-IID Data.
@@
 The highest error rate improvement for the sign-flipping attack is $0.547$, with an average error rate increase of $0.282$.
-For the Trim attack and Krum attack, the highest error rate rises are $0.451$ and $0.546$, with average error rate rises of $0.312$ and $0.278$, respectively.
+For the Trim attack and Krum attack, the highest error rate rises are $0.451$ and $0.546$, with an average error rate rise of $0.312$ and $0.278$, respectively.
 For the Min-Max attack, the highest error rate increase reaches $0.479$, with an average increase of $0.201$.
@@
 The average error rates of the global model in the cross-device FL system are lower than the error rates in the cross-silo FL system within $0.030$, illustrating our method’s generality over different data distribution and FL systems.
-Appendix G Additional Results with Fewer Malicious Clients
-We also test the error rate improvement of our method in both the IID and non-IID cross-device FL systems, with only $10\%$ malicious clients. The experimental results are shown in Table XI.
+Appendix M Additional Results with Fewer Malicious Clients
+We also test the error rate improvement of our method in both the IID and non-IID cross-device FL systems, with only $10\%$ malicious clients. The experimental results are shown in Table 10.
 Results on IID Data with Fewer Malicious Clients.
@@
 These results demonstrate the generality and robustness of our method across different data distributions and client selection methods with only a small portion of malicious clients.
+Appendix N Impact of The Pill Search Algorithm
+Figure 8: Comparison of error rates between original poisoning attacks and attacks enhanced by our method using two different pill search methods.
+We conduct a final evaluation to assess the importance and effectiveness of the ”approximate max pill search” algorithm used in our method.
+This is contrasted against a newly devised ”approximate min pill search” algorithm, which targets the least important parameters within the target model.
+Figure 8 illustrates the error rates achieved by the ”approximate max pill search”, the ”approximate min pill search”, and the original model poisoning attacks.
+The ”approximate max pill search” algorithm outperforms the ”approximate min pill search” in $41$ out of $56$ cases (approximately 7373%), underscoring its effectiveness in leveraging the most influential parameters to enhance attack impacts.
+Despite its lower efficacy, the ”approximate min pill search” still manages to surpass the original attacks in $41$ out of $56$ cases (approximately 7373%).
+This demonstrates the generality of our method across different pill search algorithms.
+Appendix O Our Method Against Adaptive Defense
+We develop an adaptive defense named DSTrust, which enhances the FLTrust’s mechanism.
+DSTrust incorporates both distance and cosine similarity scores into a unified trust score calculation, directly countering our method’s two-step adjustment approach.
+The round-$t$ trust score of client $i$ in DSTrust is calculated as follows:
+[…]
+where $\Delta\boldsymbol{g}_{t}^{(i)}$ represents the model update from client $i$ and $\Delta\boldsymbol{g}_{t}^{s})$ represents server’s model update.
+By integrating both cosine similarity and distance metrics, DSTrust provides a more comprehensive defense approach compared with FLTrust.
+This dual consideration allows DSTrust to effectively mitigate attacks that manipulate either of these metrics to bypass defenses.
+Distribution IID Non-IID
+Attack       Poison Pillw/o      Poison Pillw/       Poison Pillw/o      Poison Pillw/
+No Attack 0.108 0.116
+[…]
+[…]
+[…]
+[…]
+Table 11: Error rates under cross-silo setting against the new adaptive defense – DSTrust – with and without our method on Fashion-MNIST dataset (20% malicious clients).
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+Sign-Flipping 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
+          w/ DistAdjust       w/o SimAdjust+ Poison Pill        0.311 0.096 0.219 0.413 0.354 0.381 0.726 0.372 0.111 0.224 0.373 0.298 0.448 0.722
+          w/o DistAdjust      w/ SimAdjust+ Poison Pill       0.826 0.110 0.118 0.142 0.898 0.464 0.695 0.849 0.117 0.133 0.161 0.870 0.491 0.634
+[…]
+Trim Attack 0.243 0.109 0.139 0.146 0.174 0.179 0.116 0.332 0.120 0.201 0.163 0.231 0.238 0.124
+          w/ DistAdjust       w/o SimAdjust+ Poison Pill        0.317 0.105 0.364 0.247 0.368 0.136 0.208 0.219 0.111 0.600 0.456 0.295 0.212 0.105
+          w/o DistAdjust      w/ SimAdjust+ Poison Pill       0.554 0.104 0.122 0.108 0.429 0.284 0.119 0.744 0.110 0.134 0.126 0.898 0.478 0.109
+[…]
+Table 12: Ablation study on the necessity of both SimAdjust and DistAdjust in our method on Fashion-MNIST dataset with 20% malicious clients in the 50-client FL system.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
+Label-Flipping 0.937 0.110 0.103 0.115 0.125 0.122 0.097 0.767 0.118 0.108 0.128 0.126 0.125 0.104
+[…]
+Table 13: Error rates of Label-flipping Attack on Fashion-MNIST under cross-silo setting with 20% malicious clients in the 50-client FL system.
+Table 11 details the error rates for four baseline FL poisoning attacks both with and without our method against the DSTrust defense on the Fashion-MNIST dataset within a 50-client FL system, where 20% of clients are malicious.
+These tests were conducted under both IID and non-IID data environments.
+DSTrust effectively neutralizes the four baseline poisoning attacks when our method is not applied, highlighting its robustness as a defense mechanism.
+Despite DSTrust’s integration of both cosine similarity and distance metrics in its defense strategy, it fails to counteract the augmented attacks when our method is employed.
+Notably, our method achieves a maximum error rate increase of $0.521$, and an average error rate increase of $0.173$ across all $8$ test scenarios.
+These results demonstrate that merely understanding the adjustment strategies of our method, and subsequently integrating corresponding defense metrics, does not fundamentally negate the effectiveness of our method.
+Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics into DSTrust, it remains insufficient to thwart the enhanced capabilities of our method.
+Distribution IID Non-IID
+Attack FAvg FLT MKr Bulyan Median Trim DnC FLD Flame FAvg FLT MKr Bulyan Median Trim DnC FLD Flame
+No Attack 0.319 0.328 0.338 0.324 0.330 0.337 0.315 0.334 0.336 0.336 0.358 0.330 0.336 0.338 0.336 0.347 0.326 0.374
+Sign-Flipping 0.897 0.335 0.336 0.329 0.353 0.386 0.341 0.316 0.367 0.894 0.301 0.361 0.317 0.365 0.433 0.317 0.315 0.378
+[…]
+Trim Attack 0.431 0.323 0.422 0.428 0.434 0.432 0.340 0.339 0.362 0.423 0.345 0.469 0.429 0.474 0.451 0.352 0.299 0.355
+[…]
+Krum Attack 0.324 0.374 0.441 0.523 0.471 0.393 0.348 0.609 0.360 0.379 0.300 0.479 0.547 0.492 0.397 0.330 0.704 0.352
+[…]
+Table 14: Error rates on CIFAR-10 under cross-silo setting with 20% malicious clients and VGG-11 Net in the 30-client FL system.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.093 0.097 0.093 0.094 0.111 0.092 0.093 0.097 0.102 0.101 0.105 0.112 0.100 0.099
+Sign-Flipping 0.893 0.157 0.092 0.106 0.110 0.108 0.100 0.963 0.122 0.104 0.103 0.118 0.114 0.106
+[…]
+Trim Attack 0.274 0.126 0.108 0.105 0.191 0.219 0.097 0.357 0.135 0.134 0.308 0.225 0.270 0.101
+[…]
+Table 15: Error rates on Fashion-MNIST under cross-silo setting with 20% malicious clients in a 100-client FL system.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
+Sign-Flipping 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
++ Neurotoxin 0.710 0.147 0.105 0.103 0.106 0.105 0.110 0.739 0.146 0.121 0.122 0.124 0.124 0.119
+[…]
+Trim Attack 0.243 0.109 0.139 0.146 0.174 0.179 0.116 0.332 0.120 0.201 0.163 0.231 0.238 0.124
++ Neurotoxin 0.135 0.109 0.113 0.106 0.126 0.119 0.108 0.168 0.187 0.144 0.197 0.138 0.144 0.121
+[…]
+Table 16: Comparison with Neurotoxin under cross-silo setting on IID Fashion-MNIST in the 50-client FL system.
+Appendix P Limitations and Future Work
+Our method significantly enhances non-state-of-the-art (non-SOTA) model poisoning attacks, enabling them to SOTA results against various prevalent defenses.
+This is accomplished through a pill-based, attack-agnostic augmentation pipeline.
+We not only demonstrate our method’s capabilities but also expose fundamental vulnerabilities within the current designs of defense mechanisms.
+For future attacks in FL, it is essential for attackers to meticulously evaluate the importance of each parameter in their implementation.
+By targeting specific subsets of parameters, attackers can devise more flexible and adaptive attacks, improving stealthiness and complicating defense efforts.
+As for future defenses, while individually checking each parameter might seem viable, its practical deployment is hindered by high overheads, making it infeasible in real-world applications.
+Thus, there is a pressing need for more sophisticated defenses that can conduct fine-grained analyses of the roles of different parameters in neural networks, while executing without imposing prohibitive computational costs.
+Appendix Q Additional Ablation Study on Pill Adjustment
+To illustrate the necessity of both the SimAdjust() and DistAdjust() used in our method, we conduct a detailed ablation study, providing the error rates of the Sign-flipping Attack and the Trim Attack with different settings of pill adjustments on the IID and non-IID Fashion-MNIST dataset within a 50-client FL system in Table 12.
+The results show that using both SimAdjust() and DistAdjust() together achieves the highest error rate in multiple cases.
+For Sign-flipping Attack, the combined adjustment method outperforms others against multi-Krum (both IID and non-IID), FLDetector (both IID and non-IID), and Bulyan (non-IID).
+Similarly, for Trim Attack, the combined approach yields the highest error rates in 8 out of 14 defenses under both IID and Non-IID settings.
+In summary, the results confirm that SimAdjust() and DistAdjust() complement each other, improving attack effectiveness while maintaining stealthiness when used together.
+Appendix R Additional Results with Label-flipping Attack
+In addition to the untargeted model poisoning attacks discussed in the main text, we evaluate our method using a data-poisoning-based targeted attack: the label-flipping attack.
+Label-flipping is a straightforward yet effective targeted attack in federated learning (FL) and is also among the least stealthy data-poisoning-based attacks.
+To make the evaluation more challenging, we configured the attacker to flip all the labels of the training data on malicious clients, making the label-flipping attack even less stealthy.
+The results are shown in Table 13, demonstrating that our method enhances the label-flipping attack to bypass six additional defenses compared to its original version.
+This illustrates the compatibility of our method for data-poisoning-based and targeted attacks.
+Appendix S Additional Results with More Complex Models
+To further demonstrate the effectiveness of our method on more complex model architectures, we test our method using VGG-11 net on the IID CIFAR-10 dataset, shown in Table 14.
+The results demonstrate that our method consistently enhances the performance of Sign-flipping Attack, Trim Attack, and Krum Attack, outperforming their original versions in 41 out of 54 cases.
+These findings illustrate the effectiveness of our method with more complex model architectures.
+Appendix T Additional Results with Larger FL Systems
+To further demonstrate the effectiveness of our method in larger FL systems, we extend our experiments on the Fashion-MNSIT dataset with 100 clients, shown in Table 15.
+The results show a similar trend as observed in the 50-client system.
+Our method enables baseline attacks to successfully bypass four additional defenses, causing over 50% additional error rates in the global model.
+These findings further validate the effectiveness and generality of our approach in larger systems, when a single malicious client has fewer data samples.
+Appendix U Comparison with Existing Attack Enhancement Method
+Considering several prior studies (bagdasaryan2020backdoor; bhagoji2019analyzing; zhang2022neurotoxin) enhancing backdoor attacks, we also adapt one recent one - Neurotoxin (zhang2022neurotoxin) - to our untargeted attacks evaluation setting.
+Table 16 illustrates the results on the IID Fashion-MNIST dataset within a 50-client FL system using Trim Attack.
+The results demonstrate that our method outperforms Neurotoxin in 24 out of 28 cases.
+This highlights that directly transferring existing methods designed for backdoor attacks may not yield consistent effectiveness when applied to untargeted attack scenarios.
+The results further validate the robustness of our approach.
+Data Distribution IID Non-IID
+Attack FedAvg FLTrust MKrum Bulyan Median Trim FLD FedAvg FLTrust MKrum Bulyan Median Trim FLD
+No Attack 0.109 0.107 0.105 0.105 0.123 0.106 0.115 0.113 0.115 0.115 0.112 0.142 0.115 0.122
+Sign-Flipping 0.935 0.123 0.098 0.101 0.106 0.101 0.103 0.903 0.114 0.103 0.107 0.124 0.111 0.116
+[…]
+Trim Attack 0.102 0.112 0.101 0.106 0.113 0.103 0.095 0.118 0.113 0.118 0.113 0.136 0.114 0.108
+[…]
+Table 17: Performance when the number of malicious clients is gradually decreasing.
+Appendix V Results with Decreasing Number of Malicious Clients
+To further demonstrate the effectiveness of our method in a more practical setting, we evaluate its performance as the number of malicious clients in the FL system gradually decreases.
+Specifically, we used the Fashion-MNIST dataset within a 50-client FL system.
+Initially, 20% of clients are malicious, and for every T/4 rounds (where T is the total number of FL communication rounds), the proportion of malicious clients reduces by 5%.
+Here is a detailed breakdown of this setup:
+  • $0\rightarrow\frac{T}{4}$: 20% clients in the FL system are malicious.
+  • $\frac{T}{4}\rightarrow\frac{T}{2}$: 15% clients in the FL system are malicious.
+  • $\frac{T}{2}\rightarrow\frac{3T}{4}$: 10% clients in the FL system are malicious.
+  • $\frac{3T}{4}\rightarrow T$: 5% clients in the FL system are malicious.
+The results are presented in Table 17, demonstrating that our method significantly enhances the error rates achieved by the original Trim Attack and Sign-flipping attack in 23 out of 28 cases, with an average error rate increase of over 50%.
+These findings illustrate the robustness and effectiveness of our method in a more practical scenario.